### v1

# BAGELS: Benchmarking the Automated Generation and Extraction of Limitations from Scholarly Text

Abstract
In scientific research, “limitations” refer to the shortcomings, constraints, or weaknesses within a study. A transparent reporting of such limitations can enhance the quality and reproducibility of research and improve public trust in science. However, authors often a) underreport them in the papers’ text and b) use hedging strategies to satisfy editorial requirements at the cost of readers’ clarity and confidence. This underreporting behavior, along with an explosion in the number of publications, has created a pressing need to automatically extract/generate such limitations from scholarly papers. In that direction, we report a complete architecture for computational analysis of research limitations. Specifically, we a) create a dataset of limitations in ACL, NeurIPS, and PeerJ papers by extracting them from papers’ text and integrating them with external reviews; b) propose methods to automatically generate them using a novel Retrieval Augmented Generation (RAG) technique; c) create a fine-grained evaluation framework for generated limitations and provide a meta-evaluation for the proposed evaluation techniques. Code and datasets are here: github | huggingface
[…]
BAGELS: Benchmarking the Automated Generation and Extraction of Limitations from Scholarly Text
Ibrahim Al Azher, Miftahul Jannat Mokarrama, Zhishuai Guo, Sagnik Ray Choudhury, Hamed Alhoori Northern Illinois University, Dekalb, IL, USA University of North Texas, Denton, TX, USA {iazher, mmokarrama1, zguo, alhoori}@niu.edu, sagnik.raychoudhury@unt.edu
1 Introduction
Figure 1: System Architecture. We build a dataset by extracting limitations from a paper’s text (when mentioned) and reviews, build a generator for papers that do not explicitly mention limitations, and develop a novel framework for reference-based evaluation of generated limitations.
In the context of a scientific article, “limitations” refer to the inherent shortcomings, constraints, or weaknesses within a study that may influence the results or limit the generalizability of its findings Ross and Bibler Zaidi (2019). These limitations can arise from various elements of the research process, including the methodology, theoretical framework, data collection, experimentation, and analysis Ioannidis (2007). Limitations explicitly acknowledged by the authors often include internal validity issues, measurement errors, potential confounding factors, and the failure to measure important variables Puhan et al. (2009).
Openly discussing limitations is crucial for several reasons. It ensures the credibility and ethical standards of a scientific inquiry Bunniss and Kelly (2010); Chasan-Taber (2014); Annesley (2010). By clearly defining the boundaries of a study, researchers provide a realistic context for their findings, supporting accurate interpretation, transferability, and reproducibility Ioannidis (2007); Eva and Lingard (2008). Acknowledging limitations fosters scientific integrity by demonstrating a commitment to honesty Žydžiūnaitė (2018). It also opens pathways for refining research methods, opening areas for future work Azher et al. (2025), and leading to innovations that help future researchers avoid repeating the same shortcomings Escande et al. (2016).
Despite the benefit of honest admission of such limitations, researchers sometimes remain reluctant to include them in the text or articulate them in detail Ioannidis (2007); Ter Riet et al. (2013). Concerns about the potential impact on publication chances and, subsequently, career progression Montori et al. (2004) can encourage this biased behavior. When forced to explicitly acknowledge limitations in the text [^1] As is the practice in the NLP and ML community in recent years., authors sometimes resort to using generic or irrelevant terms, which do not meaningfully inform the reader about the study’s real constraints Ross and Bibler Zaidi (2019). Moreover, limitations can sometimes serve as a form of “hedging”, where authors present their findings cautiously to avoid making definitive statements Hyland (1998). This approach, while safer for the authors, does not contribute to the clarity or usefulness of the research.
Failure to disclose limitations undermines the scientific process and misleads readers, reviewers, and policymakers, potentially preventing them from recognizing the constrained nature of findings and possible biases Greener (2018). NLP research in automatic extraction, generation, and evaluation of limitations from scholarly papers is limited due to the lack of standardized datasets, novel methods, and evaluation framework. This study takes a step toward bridging this gap.
The number of scientific publications has exploded recently Larsen and Von Ins (2010). That, combined with the problems of self-reporting, has necessitated a need for computational methods to study the limitations of scholarly work. In that direction, our contributions are as follows (see Figure 1):
  • We create a dataset for limitations reported in research articles. For each paper, we extract the limitations from the papers’ text and OpenReview feedback  [^2] https://openreview.net/. This creates a valuable benchmark for future studies that want to extract and analyze limitations from research papers and study their impact on scholarly progress. Notably, due to the integration of authored text and external data, this benchmark suffers less from the self-reporting bias, providing a broader and more nuanced perspective about limitations.
  • We create a novel RAG system to generate limitations. This is a useful tool for editors or researchers in the under-reporting or hedging scenario.
  • Evaluating the generated limitations is a key challenge. Many traditional metrics for measuring text generation quality depend on lexical overlap (ROUGE Lin (2004), BLEU Papineni (2001)) or semantic similarity (BERTScore Zhang et al. (2019), MoverScore Zhao et al. (2019)). They are not sufficient for our purpose as they overprioritize common terms (e.g., tokens such as “bias”, “dataset”, and “generalizability” are standard in the limitations section of most AI/ML/NLP papers and should not affect the evaluation score). We develop a new evaluation framework that uses LLMs-as-judges to provide fine-grained comparisons and detailed error analysis, thus offering a more transparent and actionable evaluation for limitations.
2 Related Work
Various studies have studied the reporting, or lack thereof, of ‘limitations’ in scholarly papers. Ioannidis (2007) found that only 17% of top-tier journal articles mentioned limitations, with just 1% in abstracts. Similarly, Puhan et al. Puhan et al. (2012) identified 27% of biomedical papers lacked limitations, risking overestimated research reliability. Moreover, Goodman Goodman et al. (1994) found that in the peer review process acknowledging limitations is often problematic. Few journals require discussing limitations Ioannidis (2007), risking biased reviews and weakening scientific dialogue Horton (2002), necessitating reforms for transparency.
These gaps underscore the need for a systematic framework for limitations generation, an area that remains underexplored. Faizullah et al. Faizullah et al. (2024) propose an LLM-chain pipeline to summarize and iteratively refine candidate limitations; Al Azher et al. (2024) integrates topic modeling with LLMs to derive and elaborate structured limitation themes; and Azher (2024) develops a graph-augmented LLM approach for generating comprehensive limitation statements. Other work focuses on overcoming limitations of charts and graph by generating meanigful captions Al Azher and Alhoori (2024). However, these studies are confined to ACL/EMNLP corpora and author-stated limitations, overlooking reviewer comments (e.g., OpenReview), and evaluate only with ROUGE, BERTScore, and coherence metrics, which do not assess finer-grained topic or contextual alignment. Our proposed framework addresses existing limitations by incorporating NeurIPS papers alongside ACL, collecting OpenReview comments for each NeurIPS paper to establish a broader ground truth, and implementing a limitation-level text evaluation method to preserve granularity.
Evaluating NLP outputs is crucial for assessing quality, accuracy, and relevance. Traditional metrics like ROUGE and BLEU struggle with semantics, while BERTScore improves similarity but relies on references and lacks meaningful error analysis. Recent advancements in large language models (LLMs) have revolutionized text evaluation methods Zheng et al. (2023) using zero-shot and in-context learning Wei et al. (2022). For example, GPTScore Fu et al. (2023) uses LLMs to evaluate how likely a text could be generated based on a given context, TIGERScore Jiang et al. (2023) provide explainable error analysis, and PandaLM Wang et al. (2023) distinguish superior models among multiple models. Other approaches, such as AttrScore Yue et al. (2023), evaluate whether a reference supports or contradicts generated statements. SummacConv Laban et al. (2022) breaks documents into sentence-level units and removes summary sentences with low entailment probability.
Despite their promise, LLM-based evaluations have issues such as positioning bias, where altering the order of inputs affects results. To mitigate positioning bias, we systematically alter the position of limitations and select outputs that remain consistent across different orderings. Moreover, while these text evaluation methods focus on lexical overlap and surface-level LLM-based analysis, our proposed approach advances beyond these limitations by uniquely combining granularity-aware evaluation, considering LLM as a judge, quantitative scoring, with topic-level agreement.
3 Limitation Extraction
Our limitation extraction is twofold. At first, we extracted limitations with a science parse  [^3] https://github.com/allenai/science-parse tool. If a paper has a separate limitation section, we were able to extract it using this tool. But if the limitation has no separate section, we used Python regex to check the word ‘limitation’ and extracted all consecutive texts before another section appeared. However, this approach also collected some noisy sentences unrelated to the limitation. So, in the later part, we applied LLM and sent the texts to extract only limitations. We also used Selenium to collect reviews from OpenReview. Later, the ‘author mentioned limitation’ is sent to the zero-shot LLM to extract limitations, and the review from ’OpenReview’ is sent to the zero-shot LLM to extract shortcomings/weaknesses. Through this process, we removed the noisy sentences. (§ 3.1)
3.1 Dataset of Extracted Limitations
Granularity. A key challenge in constructing a dataset of research limitations is defining the appropriate level of granularity. Should a limitation be captured as a single phrase, a full sentence, or an entire paragraph? We define a limitation as a sequence of sentences, as individual sentences often do not encapsulate multiple limitations. In contrast, a single limitation can extend across multiple sentences, sometimes forming a complete paragraph.
Extraction Sources. Two primary sources form the basis of our dataset: (1) limitations explicitly acknowledged by authors, and (2) those highlighted through peer-review commentary. Although author-reported limitations often provide well-structured insights, previous research indicates that such limitations may be underreported or carefully hedged. To address this gap, we incorporate OpenReview [^4] https://openreview.net/ comments, in which peer reviewers frequently identify additional constraints or weaknesses not mentioned by the authors.
Our dataset consists of papers from conferences in natural language processing and machine learning: ACL [^5] https://aclrollingreview.org/cfp, NeurIPS [^6] https://neurips.cc/public/guides/PaperChecklist, and PeerJ  [^7] https://peerj.com/benefits/indexing-and-impact-factor/. We include 6,932 NeurIPS papers from 2021 and 2022 and 5,739 ACL papers from 2023-2024, 1000 biological and medical science papers from PeerJ. In addition, we integrate OpenReview comments for 2,802 papers from NeurIPS, thus capturing both self-reported limitations and those externally identified by reviewers. All of the PeerJ paper contains self-reported Limitations alongside other sections and peer review comments. The first experiment evaluates state-of-the-art LLMs’ ability to extract limitations from papers (when they are explicitly mentioned) and their reviews. For each paper, we use GPT-4o mini to extract and get an average of $8$ limitations from a paper and $10$ from open review.
LLM as a Ground Truth Limitation Extractor. Identifying and isolating limitations can be challenging, as they often appear in sections such as Conclusion, Discussion, or Future Work and may also be dispersed throughout OpenReview feedback. To address this, we implement a targeted extraction process designed to be both efficient and accurate. Rather than processing entire papers, which would be computationally expensive due to larger context windows, we first apply Python regex to identify and extract relevant sections. Some research papers include a dedicated limitations section; we extract the text of this section using the AllenAI Science Parse tool [^8] https://github.com/allenai/science-parse. When a paper doesn’t have such a well-defined section, we identify sentences containing the word limitations and continue extracting subsequent sentences until we encounter one of the following sections: Conclusion, Future Work, Ethics, Acknowledgement, or Grant. Then, we extracted the limitations from the research paper by LLM to remove noisy sentences and make segments.
To capture broader perspectives from peer reviews, we aggregate comments from multiple OpenReview responses into a single text. Finally, we employ GPT-4o  [^9] https://openai.com/index/hello-gpt-4o/ (Figure 4) to segment the extracted text and identify distinct limitation statements, distinguishing those reported by authors from those highlighted by reviewers. We evaluate the quality of these extracted limitations through a user study (§5).
Dataset Applications. The resulting dataset will be made publicly available and can be used as a benchmark for evaluating automated limitation extraction and generation methods (§5). Beyond this, the extracted limitations can be examined and organized into a taxonomy of limitations in ML and NLP, offering a more structured understanding of common research challenges. By integrating this taxonomy into citation networks, we can introduce the concept of a Limitation Multigraph, enabling scientometric analyses into whether certain limitations shape the direction of subsequent research or, alternatively, tend to be overlooked. These avenues present new opportunities to study how the reporting (or lack thereof) of limitations affects the broader scientific discourse, a topic we plan to explore in future work.
Figure 2: Evaluation of generated limitations.
4 Limitation Generation
We generated limitations by feeding every paper’s text, excluding its actual “Limitations” section into both vanilla and RAG-augmented LLMs, experimenting with either the top three sections (Abstract, Introduction, Conclusion) or the full paper as input and using three ground-truth definitions (author-stated limitations, OpenReview comments, or their combination). We computed each section’s cosine similarity to the “Limitations” section across ACL, NeurIPS dataset and found that Abstract, Introduction, and Conclusion rank highest on average (Table 19, Appendix). When full paper inputs exceed the context length, we make them chunkwise.
4.1 LLM and RAG for Limitation Generation
Most research papers do not explicitly mention limitations, or they underreport these shortcomings even when a dedicated section is provided. To address this gap, we propose a system that generates limitations directly from the paper text.
Vanilla LLM. In the vanilla LLM setup, the language model is prompted with the entire paper $P_{i}$. When the paper exceeds the context window, it is divided into chunks $\{P_{i}^{\prime}\}$, and limitations are generated for each chunk D’Arcy et al. (2024). The LLM then aggregates these chunk-specific outputs into a cohesive, meaningful final set of limitations, ensuring that no relevant material is overlooked in especially long or detailed papers.
RAG Integration. A paper $P_{i}$ can be used independently to generate limitations, but this approach risks overlooking valuable insights from other, potentially related papers. In particular, even when a paper lacks an explicit limitations section, other papers with similar methodologies or datasets may discuss relevant shortcomings. More importantly, certain findings may be implicitly contradicted by subsequent research. Also, failing to consult the broader literature leads to a significant gap in the reported limitations. To address this issue, we employ a RAG framework, which allows the system to draw on context from multiple papers rather than relying only on $P_{i}$. Specifically, we begin by splitting 100 random papers into multiple chunks and inserting these chunks into a vector database. A retriever then identifies the top $k$ chunks most relevant to the input text $\{P_{i}/P_{i}^{\prime}\}$, depending on whether the paper $P_{i}$ or a chunk $P^{\prime}_{i}$ can fit into the context window, using cosine similarity to match these chunks to the query (Figure 1, step 4,5).
5 Evaluation of Generated Limitations
We introduce a framework, called PointWise (PW), for evaluating the performance of LLMs in generating limitations for scholarly papers (Figure 2).
Problem Setup.
Suppose we have a set of papers $P=\{P_{1},P_{2},\dots,P_{n}\}$. For each paper $P_{i}$, we assume access to: Ground truth limitations $G_{i}=\{g_{i1},g_{i2},\dots,g_{ix}\}$, where $x$ is the number of ground truth limitations we extracted or annotated for $P_{i}$. And LLM-generated limitations $H_{i}=\{h_{i1},h_{i2},\dots,h_{iy}\}$, where $y$ is the number of limitations produced by the LLM for $P_{i}$. Our goal is to measure (1) how many ground truth limitations the LLM correctly reproduces (coverage) and (2) how well each matched pair of limitations aligns in content and focus (performance).
5.1 Coverage
A. Pairwise Matching.
To quantify coverage, we first create all possible pairs of limitations between the sets $G_{i}$ and $H_{i}$. Let
[…]
Hence, $|S_{i}|=x\times y$. We then use an LLM as a judge Zheng et al. (2023) to decide if a ground truth limitation $g_{ik}$ and a generated limitation $h_{il}$ are similar in content or topic:
$J(g_{ik},h_{il})\;=\;\begin{cases}1,&\text{if }g_{ik}\text{ and }h_{il}\text{ are similar},\\ 0,&\text{otherwise}.\end{cases}$
We collect all matched pairs into a set
$M_{i}\;=\;\{(g_{ik},\,h_{il})\mid J(g_{ik},\,h_{il})=1\},$
and let $\lvert M_{i}\rvert=z_{i}$ be the number of matched pairs for paper $P_{i}$.
B. Coverage of Ground Truth Limitations.
We define $C_{Gi}(g_{ik})=1$ if the ground truth limitation $g_{ik}$ appears in at least one matched pair in $M_{i}$, and 0 otherwise:
$C_{Gi}(g_{ik})\;=\;\begin{cases}1,&\exists\,h_{il}\text{ such that }(g_{ik},\,h_{il})\in M_{i},\\ 0,&\text{otherwise}.\end{cases}$
The coverage of ground truth limitations for paper $P_{i}$ is
$A_{Gi}\;=\;\frac{1}{x}\sum_{k=1}^{x}C_{Gi}(g_{ik}).$
In other words, $A_{Gi}$ measures the fraction of ground truth limitations in $P_{i}$ that are matched with at least one LLM-generated limitation.
C. Coverage of LLM-Generated Limitations.
Similarly, we define $C_{Hi}(h_{il})=1$ if a generated limitation $h_{il}$ appears in at least one matched pair in $M_{i}$, and 0 otherwise:
$C_{Hi}(h_{il})\;=\;\begin{cases}1,&\exists\,g_{ik}\text{ such that }(g_{ik},\,h_{il})\in M_{i},\\ 0,&\text{otherwise}.\end{cases}$
The coverage of LLM-generated limitations for paper $P_{i}$ is
$A_{Hi}\;=\;\frac{1}{y}\sum_{l=1}^{y}C_{Hi}(h_{il}).$
We aggregate these coverage values across all papers by taking their means:
$A_{G}\;=\;\frac{1}{n}\sum_{i=1}^{n}A_{Gi},\quad A_{H}\;=\;\frac{1}{n}\sum_{i=1}^{n}A_{Hi}.$
D. Precision, Recall, and F1.
We also compute overall precision, recall, and F1 scores. For each paper $P_{i}$:
$\text{TP}_{i}$ $=|M_{i}|,\quad\text{FP}_{i}$ $=x-\sum_{k=1}^{x}C_{Gi}(g_{ik}),$
$\text{FN}_{i}$ $=y-\sum_{l=1}^{y}C_{Hi}(h_{il}).$
Here, $\text{TP}_{i}$ (true positives) is the total number of matched pairs; $\text{FP}_{i}$ (false positives) is the number of ground truth limitations not matched by any LLM-generated limitation; $\text{FN}_{i}$ (false negatives) is the number of LLM-generated limitations unmatched by any ground truth limitation. True negatives ($\text{TN}_{i}$) are not defined since no “negative” class of limitations is specified. If there is one ground truth limitation $g_{i1}$ matches with multiple LLM-generated limitations (and vice versa), True Positive (TP) counts as one. (Details in Appendix A.1)
5.2 Performance
After finding matched pairs $(g_{ik},h_{il})\in M_{i}$, we evaluate their quality using text based evaluation (Rouge-1, Rouge-2, BERTSCore, Cosine Similarity), keyword based evaluation, and heading based evaluation (Details in Appendix A.2).
6 Experimental Setup
We fine-tuned three sequence-to-sequence models, T5 (512-token window) Raffel et al. (2020), BART (1024 tokens), and Pegasus (1024 tokens) Zhang et al. (2020) on a 70 / 30 train–test split. All were trained for 3 epochs with a learning rate of $5\times 10^{-5}$, weight decay 0.01, 300 warmup steps, and batch sizes of 4 (train) and 8 (eval), with early stopping; inputs longer than 512 tokens were truncated. For zero-shot evaluation, we used, GPT-3.5 (16 K) and GPT-4o (128 K) at temperature 0.1, and Llama 3 8B and Mistral 7B (128 K) via Ollama. To incorporate retrieval (RAG), we built a VectorStoreIndex with llama-index  [^10] https://www.llamaindex.ai/., and OpenAI text-embedding-ada-002 embedding model for encoding the source and query documents. We took 100 random papers from the dataset and stored them in a vector database. These documents were chunked into 800-token segments with 20-token overlaps to preserve context, and the LLM temperature was set to 0.
7 Experiments and Results
7.1 Limitation Extraction Evaluation
We manually collected limitation sections from 500 ACL papers randomly (human curated). We measure the n-gram overlap and semantic similarity between human-curated limitations vs. automatically extracted ground truth limitations to evaluate our automatic limitation extraction process. As shown in Table 1, our automatically extracted limitation text achieves strong alignment with human-curated ground truth limitations on the ACL ’23 and ACL ’24 datasets, with ROUGE-1, L scores exceeding 85% and BERTScore above 95%.
Metrics ACL 24 ACL 23
ROUGE-1 86.68 85.07
ROUGE-2 83.13 81.68
ROUGE-L 86.64 85.04
BLEU 77.70 75.15
Jaccard Similarity 79.36 77.08
JS Divergence 23.58 24.88
Levenshtein 92.54 90.32
Cosine Similarity 90.05 88.70
BERTScore (F1) 95.70 95.66
Table 1: Evaluating the Alignment of Human-Collected vs. Automated Limitation Extraction.
Human Evaluation. We employ three annotators (separate from this paper’s authors) [^11] CS graduate students with research experience in NLP and AI to evaluate the quality of the extracted limitations. We choose a sample of 100 limitations, and for each, we show them the source and ask a Yes/No question, whether they thought the LLM extracted the limitation from the source, as opposed to generating it from scratch. Each annotator answer positively in $>90\%$ of cases (avg $\pm$ std= $95\pm 2.45\%$) (Table 2). The annotator pairs have weighted Kappa scores between $38-75\%$, denoting fair to substantial agreement (Table 18, Appendix). We checked Llama 3 70b as an extractor, but the extraction quality was not good. Overall, the extraction quality is high, and the annotators agree; therefore, we use the extracted limitations as ground truth in the subsequent evaluations. Furthermore, two annotators manually verified the extracted limitations from 1000 papers. They assessed whether each automatically extracted limitation (via tool + LLM) was grounded in the original paper text or OpenReview content and whether it included any irrelevant or unrelated information. Their analysis confirmed that all extracted limitations were faithfully sourced, with no instances of hallucinated, noisy, or newly generated content.
Model Role Sample U1 U2 U3
GPT 4o mini Extractor 100 92 95 98
Table 2: Evaluating LLM as an extractor role with human annotator (U).
Model R-L BS JStext CStext CGT CLLM Prec. Recall F1
T5 19.92 87.81 10.82 31.79 35.48 29.59 0.29 0.31 0.30
BART 19.43 87.67 10.68 31.91 33.71 30.10 0.30 0.31 0.31
Pegasus 20.15 87.66 10.71 33.39 29.28 25.27 0.25 0.26 0.26
Llama 3 25.66 88.30 14.69 40.4 61.38 39.04 0.39 0.50 0.44
GPT 3.5 24.24 87.08 14.65 43.12 76.62 46.65 0.47 0.67 0.55
GPT 3.5 + RAG 30.21 90.88 19.47 45.37 39.99 44.66 0.42 0.40 0.41
GPT 4o-mini 16.57 86.02 8.70 32.29 57.65 19.76 0.20 0.31 0.24
GPT 4o mini + RAG 23.17 87.33 12.99 39.29 67.13 45.67 0.57 0.45 0.51
Table 3: Results of models in “Coverage” (Coverage of Ground Truth Limitation (CGT), LLM Generated Limitation (CLLM), Precision, Recall, and F1-score) and “performance” metrics – Rouge-x, BLEU, BertScore (BS), Jaccard (JS) and Cosine (CS) similarity over text (ACL data).
7.2 Limitation Generation Evaluation
Coverage Evaluation. Table 3 reports results on ACL data using only author-stated limitations as ground truth. We couldn’t collect the OpenReview of ACL papers, so we used the author-mentioned limitations as ground truth. GPT-3.5 achieves the highest Coverage of Ground Truth (CGT=76.62) and Coverage of LLM-Generated Limitations (CLLM=46.65). Vanilla LLMs tend to produce overly verbose limitations. Integrating RAG yields more focused, concise outputs that cover a broader range of topics and align better with ground truth—for both GPT-3.5 and GPT-4o mini. In particular, incorporating RAG increases CLLM by 25.91 points over GPT-4o mini. For NeurIPS data (Table 4), considering RAG increase CGT, decrease CLLM, which gains F1 score (+0.06). Compared to ACL, NeurIPS, and Peerj, the ACL dataset archives the highest CGT, and NeurIPS achieves the highest CLLM (Table 11, Appendix).
Performance Evaluation. Table 3 reports our performance metrics using author-extracted limitations as ground truth. Compared to all models, GPT 3.5 + RAG shows the best results. Augmenting GPT-3.5 with RAG boosts every measure, ROUGE-L (+5.97) and BERTScore (+3.8). GPT-4o mini exhibits the same upward trend when paired with RAG. For NeurIPS data (Table 4), considering RAG increased the performance of all metrics except Cosine Similarity. Among our datasets, NeurIPS achieves the highest ROUGE, BERTScore, and BLEU largely due to the inclusion of OpenReview feedback (Table 11, Appendix). Moreover, our proposed PointWise approach achieves better BERTScore than the traditional way in every model (Table 8, 17, Appendix).
Metric GPT 4o mini GPT 4o mini + RAG
Rouge-L 17.07 18.10 (+1.03)
BERTScore 86.25 86.45 (+0.2)
JStext 9.12 9.72 (+0.6)
CStext 35.42 34.71 (-0.71)
CGT 45.69 61.76 (+16.07)
CLLM 76.27 59.18 (-17.09)
F1 score 0.54 0.60 (+0.06)
Table 4: Coverage and Performance measure using GPT 4o mini + RAG in NeurIPS data. Ground truth is Limitation + OpenReview, and input is the top 3 sections alongside the abstract.
Evaluation ‘LLM as a Judge’. We evaluated our Judge LLM model GPT 4o mini regarding PointWise evaluation approach. From each pair of ground truth and LLM generated limitation, we asked LLM whether both are same or not (Figure 3). We use the same annotators as before and choose a sample of 200 pairs, 100 positive cases where the LLM predicted a match and 100 negatives. We ask humans the same question as the model and calculate the weighted Kappa scores. As seen in (Table 9, Appendix), the human experts overwhelmingly agree with GPT, with kappa values ranging between $90$-$95\%$. The agreement between human annotators is $>92\%$, indicating a perfect agreement. We also experimented with a strong baseline of Llama-3.1 400B, which shows a poor agreement with humans.
7.3 Ablation Study
We ablate in two dimensions a) the size of the input text. b) the ground truth
a. Size of the input text: Table 5 presents results using GPT-4o mini. Moving from the top three sections to all sections as input improves every metric across all ground-truth types. The only exceptions are a few cases, such as BERTScore and LLM Coverage (CLLM) when using OpenReview as a ground truth. We also experimented ablation study with other models (Table 6), such as Mistral 7B and Llama 3 8B on NeurIPS data. Using all sections as input yields more detailed outputs, boosting LLM coverage by 8.35 points. However, it slightly reduces ground-truth coverage, CGT (-0.29), ROUGE-L (+0.64), and BERTScore (+0.36). Llama 3 8B shows the opposite trend: using all sections significantly degrades its performance. This likely reflects Llama’s limited context window, which hinders its ability to process longer inputs and generate coherent limitations.
b. Ground truth: Table 5 compares zero-shot GPT-4o mini results against three ground-truth sets: author-stated limitations, OpenReview comments, and their combination. Using both Limitations + OpenReview yields higher scores on nearly every metric compared to Limitations alone, except for CSKeyword, JSKeyword, and CGT. This demonstrates that reviewer feedback enriches the ground truth, improving overall performance, but because the combined ground truth is larger, its relative coverage (CGT) can decrease. We also experimented with Mistral 7b (Table 7, Appendix), incorporating OpenReivew increased CLLM, Cosine similarity, and F1 score but marginal drops in other metrics.
Metric Input Ground Truth
Sec Auth OPR Auth + OPR
R-L 3 16.57 18.55 17.07
(+0.5)
R-L All 17.11 (
$\uparrow$
18.65 (
$\uparrow$
18.83
(+1.72)
BS 3 86.02 86.50 86.25
(+0.23)
BS All 86.24 (
$\uparrow$
85.96 (
$\downarrow$
86.67
(+0.43)
JStext 3 8.70 10.17 9.12
(+0.42)
JStext All 9.22 (
$\uparrow$
10.23 (
$\uparrow$
10.38
(+1.16)
CStext 3 32.29 38.35 35.42
(+3.13)
CStext All 33.23 (
$\downarrow$
38.74 (
38.87
(+5.64)
CGT 3 57.65 54.79 45.69
(-12.41)
CGT All 61.55 (
$\uparrow$
56.55 (
$\uparrow$
49.43
(-12.12)
CLLM 3 19.76 56.79 76.27
(+57.07)
CLLM All 40.04 (
$\uparrow$
56.77 (
$\downarrow$
42.59
(+2.55)
F1 3 0.24 0.55 0.54
(+0.3)
F1 All 0.45 (
$\uparrow$
0.56 (
$\uparrow$
0.44
(-0.01)
Table 5: GPT 4o mini results in “coverage” (Coverage of Ground Truth Limitation (CGT), LLM Generated Limitation (CLLM), F1-score) and “performance” (Rouge-x, BertScore (over entire text and just heading), Jaccard (JS) and Cosine (CS) similarity over text and heading) (NeurIPS data).
Metric Input (3 Sec.) Input (All Sec.)
Mistral 7B
Rouge-L 14.59 13.95 (-0.64)
BERTScore 84.59 84.23 (−0.36)
JStext 7.70 7.49 (−0.21)
CStext 30.95 31.82 (+0.87)
BLEU 0.30 0.30 (0)
CGT 38.28 37.99 (−0.29)
CLLM 22.09 30.44 (+8.35)
F1 score 0.24 0.32 (+0.08)
Llama 3 8B
Rouge-L 25.66 22.78 (−2.88)
BERTScore 88.30 87.29 (−1.01)
JStext 14.69 11.16 (−3.53)
CStext 40.40 34.80 (−5.60)
JSkeyword 20.28 20.31 (+0.03)
CGT 61.38 18.01 (−43.37)
CLLM 39.04 5.01 (−34.03)
F1 score 0.44 0.05 (−0.39)
Table 6: Ablation study comparing Mistral 7B and Llama 3 8B on NeurIPS data. We report metrics with only the top 3 sections versus all sections as input, using author-mentioned limitations as ground truth.
8 Discussion
Our results demonstrate that leveraging external OpenReview feedback alongside author-stated limitations substantially enriches the scope and quality of LLM-generated limitations, improving performance and coverage. This indicates the model not only reproduces author-reported issues but also uncovers additional shortcomings highlighted by peer reviewers. Similarly, integrating a Retrieval-Augmented Generation (RAG) module with GPT-3.5/4 improved every metric by grounding the LLM in relevant external context. The RAG-enhanced model produced more concise, targeted limitations (fewer sentences on average), trading a slight dip in recall with improved performance. Using all sections marginally improved performance then considering 3 section in most of the metrics. Moreover, our PointWise (PW) framework breaks each limitation into fine-grained points. It employs an LLM "judge" to evaluate alignment one by one, where traditional lexical metrics can’t capture depth and relevance. Finally, chunking long documents ensures no critical content is lost outside an LLM’s context window, and RAG helps focus on significant, relevant limitations outside the current paper. Moreover, models like GPT-4o mini deliver consistently strong performance with high correlation with humans as extractors and judges, whereas others (e.g., Llama) occasionally introduce noise. These findings affirm that combining external review data, targeted section selection, RAG augmentation, and a detailed PointWise evaluation yields more trustworthy and comprehensive LLM-generated limitations—and paves the way for future work in automated scientific limitations.
9 Conclusion
We present a new approach for automatically extracting, generating and evaluating limitations in scientific articles. Our method explores chunk analysis, accommodating top sections of the entire paper and integrating OpenReview feedback to capture perspectives beyond those of the original authors. To evaluate the effectiveness of our system, we introduce a granular text evaluation framework that breaks down limitations into more minor points and employs LLMs as a Judge for assessing topic-level alignment. Combining standard NLP metrics with a keyword-based analysis provides a more nuanced and reliable assessment. Our findings are further validated by human review, which reveals a strong alignment between human judgments and the LLM-driven extraction process, including when the LLM is used as a judge.
Limitations
In this work, we focused on venues in natural language processing (ACL papers from 2023-2024) and machine learning (NeurIPS papers 2021-2022), and Biology domain papers from PeerJ, which ensures high relevance and quality but insufficient for broader generalizability. While this scope allows us to benchmark the performance of LLMs in extracting limitations from well-structured scientific texts, we acknowledge that the findings may not generalize to papers from other fields, such as social sciences, physics, chemistry, or mathematics where writing conventions and limitation styles may differ. Due to high API costs, we did not experiment with GPT-4 or GPT-4o; instead, we opted for GPT-4o Mini as a cost-effective alternative. Similarly, to avoid the high expense of maintaining an extensive vector database for RAG with closed LLMs, we limited our retrieval corpus to 100 randomly selected papers. We did not evaluate Open LLMs with RAG, such as Llama and Mistral, to assess their performance in retrieval-augmented settings. While we incorporated OpenReview comments for NeurIPS papers, we could not find them for ACL papers. Furthermore, we relied only on GPT-4o Mini as the evaluation judge and did not experiment with other LLMs for assessment. To evaluate the effectiveness of LLMs as both text extractors and judges, we conducted a human annotation study with 200 samples and only three annotators.
We used GPT 4o mini for text extraction, text generation, and LLM as a judge. We experimented with other LLMs such as Llama but GPT 4o mini performs better. To rely on LLM for limitation extraction and evaluation ther are some potential biases.
  • Self-Validation Bias: We used the same LLM for text generation and evaluation in experiments. Where a model can favor its style or logic.
  • Positional bias: For LLM as a judge, we have multiple texts to check whether both are similar. Here, LLM can have positional bias, which is more centered on the first input text.
  • Confirmation bias: We experimented with the same LLM (GPT 4o mini) for text extraction and generation. Even though both are used in an isolated environment with zero shot (no fine-tuning), the same LLM can produce confirmation bias on some common/generic limitations such as “Generalizability issue" or “small sample size" which reduces novelty.
  • Hallucination and Overconfidence: When extracting limitations, LLM might sometimes invent limitations that aren’t grounded in the source text.
To mitigate self-validation bias, we compared our LLM’s judgments against human evaluation, tested alternative models for generation and judgment (Llama and Mistral), and incorporated RAG. We changed the order and considered the consistent one to overcome the positional bias problem. Moreover, we integrated RAG with LLM for text generation, where LLM can get diverse evidence or force diversity sampling on generations to minimize the confirmation bias problem. To overcome hallucination and overconfidence, we evaluate the extraction process by LLM with three annotators to determine whether LLM extracts limitations without inventing anything new. However, these approaches are not sufficient.
As part of our future work, we plan to expand the dataset to include papers from a broader range of domains, such as arXiv submissions from diverse subject areas (e.g., bioinformatics, cognitive science) and other peer-reviewed venues beyond ACL and NeurIPS. We also aim to incorporate more review texts from platforms, where available, across different conferences. This expansion will allow us to test the robustness of our limitation extraction models across domains and improve generalizability. We will explore broader contexts from open LLMs incorporating RAG, and integrating more extensive multi-agent LLM frameworks. Moreover, we will gather annotations from a larger and more diverse pool of experts, enabling a deeper analysis of generated limitations. In doing so, we aim to make our framework more robust and broadly applicable while maintaining its core strength of providing granular, interpretable assessments of LLM-generated limitations.
Ethics Statement
This research adheres to the ethical standards in scientific research. All data used in this study, including research papers and OpenReview feedback, were obtained from publicly available sources and used in compliance with their respective usage policies. No private or sensitive information was accessed or utilized. We acknowledge the potential biases inherent in large language models (LLMs) and have taken steps to mitigate their impact. For instance, we conducted human evaluations to validate the outputs of LLMs, ensuring alignment with human judgment and reducing the risk of generating misleading or inaccurate content. Additionally, our evaluation framework emphasizes transparency and granularity, enabling a more nuanced assessment of generated limitations. For the user study, we engaged three computer science graduate students; they are no longer involved in this research. Their participation was voluntary, and they had no conflicts of interest regarding the study. We are committed to promoting the responsible use of AI in scientific research. Our work aims to enhance scientific studies’ transparency, reliability, and reproducibility by generating and evaluating limitations systematically and ethically. We encourage further research to address potential biases in LLMs and improve their application in scientific contexts. Our study didn’t filter out any dataset based on race, ethnicity, color, location, gender, or other discriminatory attributes. Therefore, we are dedicated to complying with the ethical code and policy attributed to ACL. Large language models inherently carry biases and risks derived from their training corpora—confirmation bias toward “safe” limitations (e.g., small sample size), fluency bias favoring well-written over factually grounded outputs, verbosity bias rewarding longer responses, and self-validation bias when the same model is used for extraction, generation, and evaluation. To mitigate these, we ground all generations in source content and peer reviews via Retrieval-Augmented Generation (RAG), diversify our ground truth by incorporating human-authored OpenReview critiques, and conduct parallel human evaluations to detect overconfidence and evaluation bias. RAG-based outputs are demonstrably more concise, reducing verbosity bias, and our multi-model judge setup (including Llama and Mistral alongside GPT-4o mini) helps break circularity. We acknowledge that further work is needed to quantify these biases rigorously and plan to investigate cross-domain robustness and fairness in future studies.
Appendix A Appendix
In our PointWise evaluation method, we measured precision, recall, and F1 score from True Positive, False Positive, and False Negative.
A.1 Coverage Measurement
We compute:
$\text{P}_{{r}_{i}}\;=\;\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FP}_{i}},\quad\text{R}_{{r}_{i}}\;=\;\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FN}_{i}},$
and the F1 score is the harmonic mean of $\text{P}_{{r}_{i}}$ and $\text{R}_{{r}_{i}}$.
A.2 Performance Measurement
A. Text-Based Evaluation.
We apply standard text similarity metrics to each matched pair, including ROUGE-1, ROUGE-L, BERTScore, Cosine Similarity, Jaccard Similarity, and BLEU, calculating the number of overlapping unigrams, the longest sequence of words, and the similarity between contextual embeddings.
B. Keyword-Based Evaluation.
We employ KeyBERT Grootendorst (2020) to extract a set of top keywords from the ground truth limitations $K_{G_{i}}$ and from the LLM-generated limitations $K_{H_{i}}$. We then measure the cosine and Jaccard similarity between $K_{G_{i}}$ and $K_{H_{i}}$ for each paper $P_{i}$ and average these scores across the dataset.
C. Heading-Based Evaluation.
We also compare concise “headings” or short titles for each limitation. Let $T_{G_{i}}$ be the heading for $G_{i}$ and $T_{H_{i}}$ the heading for $H_{i}$. We compute BERTScore between $T_{G_{i}}$ and $T_{H_{i}}$ for every paper $P_{i}$ and then average these values. This provides a high-level measure of how closely the top-level concepts align.
By combining coverage and performance metrics in a PointWise manner, our framework provides a detailed assessment of how well an LLM-generated set of limitations captures the breadth and depth of the ground truth. This approach also facilitates fine-grained error analysis by examining matched pairs on a per-limitation basis.
We measure coverage for both ground truth and LLM-generated limitations independently, focusing on each unique limitation within the matched pairs.
Furthermore, we conduct experiments using:
  1. 1.
    The top three sections (Abstract, Introduction, and Conclusion)
  2. 2.
    The entire paper (full paper)
This setup enables us to examine how restricting the analysis to specific sections affects coverage and matching performance.
We used three distinct prompts to check the topic level similarity between ground truth limitations and LLM-generated limitations (Figure 3, Appendix). To overcome the position bias, we choose the consistent one.
Metric Limitation Limitaiton + OPR
Rouge-L 14.59 14.52 (-0.07)
BERTScore 84.59 84.54 (-0.05)
JStext 7.70 7.62 (-0.08)
CStext 30.95 32.02 (+1.07)
CGT 38.28 32.51 (-5.77)
CLLM 22.09 44.46 (+22.37)
F1 score 0.24 0.34 (+0.1)
Table 7: Ablation study with Mistral 7b in NeurIPS data, considering Limitation and Limitation + OpenReview (OPR) as ground truth. Top 3 sections are input here.
We also compared BERTScore between the Traditional approach vs the Pointwise approach, where the Pointwise approach preserves nuanced details and showed better performance (Table 8, Appendix).
Approach Traditional way Pointwise approach
T5 82.60 87.81 (+5.21)
BART 83.09 87.67 (+ 4.58)
Pegasus 82.76 87.66 (+4.9)
GPT 3.5 83.79 87.08 (+3.29)
GPT 4o mini 82.66 86.02 (+3.36)
Table 8: Comparison of Traditional Approach and Pointwise Approach. BERTScore of each model in ACL data (Limitation is ground truth)
GPT-4 Llama HE1 HE2 HE3
GPT-4 1 0.71 0.9 0.92 0.95
Llama - 1 0.81 0.79 0.76
HE1 - - 1 0.98 0.95
HE2 - - - 1 0.97
HE3 - - - - 1
Table 9: Evaluating how good LLM ‘as a Judge’ by checking Human Expert (HE) and model (GPT-4o mini, Llama-3.1 400B) agreement in determining whether an extracted limitation matches a generated one (in PointWise Evaluation).
Model R-1 R-2 R-L BertScore BertScoreheading JStext CStext BLEU CSkeyword JSkeyword
T5 27.13 5.11 19.92 87.81 - 10.82 31.79 0.4 18.69 11.26
BART 26.05 5.49 19.43 87.67 - 10.68 31.91 0.7 19.89 12.19
Pegasus 27.0 5.46 20.15 87.66 - 10.71 33.39 0.6 17.65 10.58
GPT 3.5 32.74 10.92 24.24 87.08 85.50 14.65 43.12 3.54 26.54 17.46
Llama 3 33.98 12.02 25.66 88.30 85.30 14.69 40.4 3.6 30.33 20.28
GPT 4o-mini 22.90 3.99 16.57 86.02 75.02 8.70 32.29 0.33 16.20 9.85
GPT 3.5 + RAG 38.11 16.31 30.21 90.88 87.65 19.47 45.37 6.64 35.74 24.65
GPT 4o mini + RAG 31 9.95 23.17 87.33 - 12.99 39.29 2.69 26.47 17.19
Table 10: Results of models in “performance” metrics – Rouge-x, BLEU, BertScore (over entire text and just heading), Jaccard (JS) and Cosine (CS) similarity over text and heading (ACL data).
Metric ACL NeurIPS PeerJ
Rouge-L 16.57 17.07 10.62
BERTScore 86.02 86.25 83.26
JStext 8.70 9.12 5.81
CStext 32.29 35.42 26.19
CSkeyword 16.20 15.67 2.89
JSkeyword 9.85 9.31 1.66
BLEU 0.33 0.4 0.01
CGT 57.65 45.69 55.07
CLLM 19.76 76.27 35.87
F1 score 0.24 0.54 0.39
Table 11: Comparison among ACL and NeurIPS PeerJ data in terms of Coverage and Performance. Full text is input, using GPT 4o mini to generate limitations, and Ground truth is Limitation + Peer Review (for NeurIPS and PeerJ), ACL has only author-mentioned Limitations.
Metric Input Ground Truth
Sec Auth OPR Auth + OPR
R-1 3 22.90 26.13 23.69 ( +0.79)
R-1 All 23.74 (
$\uparrow$
26.36 (
$\uparrow$
26.36 (+2.62)
R-2 3 3.99 5.44 4.75
(+0.76)
R-2 All 4.51 (
$\uparrow$
5.41 (
$\downarrow$
5.64
(+1.13)
R-L 3 16.57 18.55 17.07
(+0.5)
R-L All 17.11 (
$\uparrow$
18.65 (
$\uparrow$
18.83
(+1.72)
BS 3 86.02 86.50 86.25
(+0.23)
BS All 86.24 (
$\uparrow$
85.96 (
$\downarrow$
86.67
(+0.43)
BSheading 3 75.02 85.79 69.50
(-5.52)
BSheading All 85.37 (
$\uparrow$
85.75 (
$\downarrow$
85.59
(+0.22)
JStext 3 8.70 10.17 9.12
(+0.42)
JStext All 9.22 (
$\uparrow$
10.23 (
$\uparrow$
10.38
(+1.16)
CStext 3 32.29 38.35 35.42
(+3.13)
CStext All 33.23 (
$\downarrow$
38.74 (
38.87
(+5.64)
BLEU 3 0.33 0.5 0.4
(+0.07)
BLEU All 0.4 (
$\uparrow$
0.5 (
$\uparrow$
0.6
(+0.2)
CSKeyword 3 16.20 15.15 15.67
(-0.53)
CSKeyword All 18.14 (
$\uparrow$
15.42 (
$\uparrow$
16.83
(-2.72)
JSKeyword 3 9.85 8.99 9.31
(-0.54)
JSKeyword All 11.10 (
$\uparrow$
9.21 (
$\uparrow$
10.19
(-0.91)
CGT 3 57.65 54.79 45.69
(-12.41)
CGT All 61.55 (
$\uparrow$
56.55 (
$\uparrow$
49.43
(-12.12)
CLLM 3 19.76 56.79 76.27
(+57.07)
CLLM All 40.04 (
$\uparrow$
56.77 (
$\downarrow$
42.59
(+2.55)
Pr. 3 0.20 0.56 0.65
(+0.46)
Pr. All 0.4 (
$\uparrow$
0.57 (
$\uparrow$
0.43
(+0.03)
Rec. 3 0.31 0.55 0.46
(+0.15)
Rec. All 0.51 (
$\uparrow$
0.56 (
$\uparrow$
0.46
(-0.05)
F1 3 0.24 0.55 0.54
(+0.3)
F1 All 0.45 (
$\uparrow$
0.56 (
$\uparrow$
0.44
(-0.01)
Table 12: GPT 4o mini results in “coverage” (Coverage of Ground Truth Limitation (CGT), LLM Generated Limitation (CLLM), Precision, Recall, and F1-score) and “performance” (Rouge-x, BertScore (over entire text and just heading), Jaccard (JS) and Cosine (CS) similarity over text and heading) metrics when different input sizes and sources are used for limitation extraction. Increasing input length and including external reviews improve generation results (NeurIPS data).
Metric GPT 4o mini GPT 4o mini + RAG
Rouge-1 23.69 24.69 (+1.0)
Rouge-2 4.75 5.24 (+0.49)
Rouge-L 17.07 18.10 (+1.03)
BERTScore 86.25 86.45 (+0.2)
BSheading 69.50 74.74 (+5.24)
JStext 9.12 9.72 (+0.6)
CStext 35.42 34.71 (-0.71)
CSkeyword 15.67 15.86 (+0.19)
JSkeyword 9.31 9.60 (+0.29)
BLEU 0.4 0.7 (+0.3)
CGT 45.69 61.76 (+16.07)
CLLM 76.27 59.18 (-17.09)
Precision 0.65 0.60 (-0.05)
Recall 0.46 0.59 (+0.13)
F1 score 0.54 0.60 (+0.06)
Table 13: Coverage and Performance measure using GPT 4o mini + RAG in NeurIPS data. Ground truth is Limitation + OpenReview, and input is the top 3 sections alongside the abstract.
Metric Input (3 sections) Input (All Sections)
Rouge-1 21.32 20.5 (-0.82)
Rouge-2 3.91 4.0 (+0.09)
Rouge-L 14.59 13.95 (-0.64)
BERTScore 84.59 84.23 (-0.36)
JStext 7.70 7.49 (-0.21)
CStext 30.95 31.82 (+0.87)
BLEU 0.3 0.3 (0)
CSkeyword 16.43 17.29 (+0.86)
JSkeyword 10.00 10.59 (+0.59)
CGT 38.28 37.99 (-0.29)
CLLM 22.09 30.44 (+8.35)
F1 score 0.24 0.32 (+0.08)
Table 14: Ablation study with Mistral 7b in Neurips data. Considering top 3 sections and all sections as input. Author mentioned limitation is ground truth here.
Metric Input (3 sections) Input (All Sections)
Rouge-1 33.98 29.57 (-4.41)
Rouge-2 12.02 9.43 (-2.59)
Rouge-L 25.66 22.78 (-2.88)
BERTScore 88.30 87.29 (-1.01)
JStext 14.69 11.16 (-3.53)
CStext 40.40 34.80 (-5.6)
BLEU 3.61 1.22 (-2.39)
CSkeyword 30.33 30.73 (+0.4)
JSkeyword 20.28 20.31 (+0.03)
CGT 61.38 18.01 (-43.37)
CLLM 39.04 5.01 (-34.03)
F1 score 0.44 0.05 (-0.39)
Table 15: Ablation study with Llama 3 8b in ACL data. Considering top 3 sections and All sections as input. Author mentioned limitation is ground truth here.
Metric Limitation Limitaiton + OpenReview
Rouge-1 21.32 21.02 (-0.3)
Rouge-2 3.91 3.74 (-0.17)
Rouge-L 14.59 14.52 (-0.07)
BERTScore 84.59 84.54 (-0.05)
JStext 7.70 7.62 (-0.08)
CStext 30.95 32.02 (+1.07)
BLEU 0.3 0.2 (-0.1)
CSkeyword 16.43 14.69 (-1.74)
JSkeyword 10.00 8.76 (-1.24)
CGT 38.28 32.51 (-5.77)
CLLM 22.09 44.46 (+22.37)
F1 score 0.24 0.34 (+0.1)
Table 16: Ablation study with Mistral 7b in NeurIPS data, considering Limitation and Limitation + OpenReview as ground truth. Top 3 sections are input here.
Approach Limitation OpenReview Limitaiton + OpenReview
Mistral (Traditional) 70.67 70.19 70.38
Mistral (Pointwise) 84.23 (+13.56) 84.59 (+14.4) 84.54 (+14.16)
Llama (Traditional) 67.55 67.17 67.50
Llama (Pointwise) 87.29 (+19.74) - -
Table 17: BERTScore Comparison Between PointWise (Proposed) and Traditional Approaches Across three different Ground Truths (Limitaiton, OpenReview, Limitation + OpenReview) in NeurIPS Data
Agreement User 1 User 2 User 3
User 1 1 0.75 0.38
User 2 - 1 0.55
User 3 - - 1
Table 18: Inter annotator agreement amont human annotator for GPT 4o mini as Extractor.
Section Cosine Similarity
Abstract vs Limitation 33.27
Introduction vs Limitation 33.06
Related Work vs Limitation 25.10
Methodology vs Limitation 26.58
Dataset vs Limitation 25.59
Conclusion vs Limitation 33.04
Experiment and Results vs Limitation 31.73
Table 19: Cosine Similarity between each section and the Limitation section.
Model Name Top 3 Sections All Sections
CGT CLLM F1 Score CGT CLLM F1 Score
Llama 3 61.38 39.04 0.44 18.01 (-43.37) 5.01 (-34.03) 0.05 (-0.39)
GPT-4o Mini 57.65 19.76 0.24 61.55 (+3.9) 40.28 (+20.52) 0.46 (+0.22)
Mistral 38.28 22.09 0.24 37.99 (-0.29) 30.44 (+8.35) 0.32 (+0.08)
Table 20: Performance (Coverage) of various models considering the top three sections versus the entire paper.
Note: CGT, CLLM denotes Coverage of Ground Truth Limitation, Coverage of LLM Generated Limitation. The ’+’ sign in parentheses indicates the performance gain achieved by using all sections instead of only the top three sections.
Prompt 1 = ''' A tuple contains (list1, list2). Check whether both ’list1’ and ’list2’ have similar topics or limitation. If both have similar topics or limitations you can say "Yes", otherwise "No". Your answer should be "Yes" or "No" with explanation. '''   Prompt 2 = ''' A tuple contains (list2, list1). Check whether both ’list2’ and ’list1’ have similar topics or limitation. If both have similar topics or limitations you can say "Yes", otherwise "No". Your answer should be "Yes" or "No" with explanation.   Prompt 3 = ''' Check whether ’list2’ contains a topic or limitation from ’list1’ or ’list1’ contains a topic or limitation from ’list2’. Your answer should be "Yes" or "No" with explanation.
Figure 3: LLM as a Judge for each limitation.
Note: We use three distinct prompts to verify consistency.
Ground Truth Limitation LLM Generated Limitation BS BS(h) Match
Error-Prone Models:
- Evaluation at the grade school mathematics level indicates that LLMs are still error-prone, but this does not provide insight into their effectiveness in tutoring other subjects.
Keywords: tutoring, LLMs, mathematics, models, subjects
Reliance on Large Language Models (LLMs): The research paper’s dependence on LLMs as Intelligent Tutoring Systems (ITS) for elementary school mathematics may limit the generalizability of the findings.
Keywords: tutoring, mathematics, elementary, models, language
87.88 81.08 Yes
Errors in Code Generation: The presence of significant errors in code generation indicates that LLMs still face challenges in accurately performing tutoring tasks.
Keywords: tutoring, LLMs, errors, code, challenges
90.05 81.38 Yes
Table 21: Ground Truth Limitation matched with multiple LLM Generated Text in PointWise Evaluation.
Note: BS, BS(h), Match denotes BERTScore, BERTSCore (heading), and LLM as a judge decision.
Prompt 1 = ''' Here is the text containing extracted limitations. Please identify and list each limitation, ensuring that each one addresses a distinct topic or point. '''
Figure 4: Prompt to extract limitations from ground truth text.
Ground Truth Limitation LLM Generated Limitation BS BS(h) Match
Quality of Representations vs. Range of Words: The primary limitation is not the quality of the word representations produced, but rather the restricted range of words that the methodologies can be applied to.
Keywords: representations, words, limitation, methodologies, range
Dictionary Limitations: Dictionaries may not include definitions for all word forms in these languages, which restricts the applicability of the proposed approach.
Keywords: dictionaries, dictionary, languages, forms, definitions
89.12 83.08 Yes
Dependence on Bilingual Dictionaries: The approach relies on bilingual dictionaries, which may not cover all necessary vocabulary or nuances of the languages.
Keywords: bilingual, dictionaries, languages, vocabulary, relies
91.71 82.78 Yes
Table 22: LLM Generated Limitation matched with multiple Ground Truth Limitation.
Note: BS, BS(h), Match denotes BERTScore, BERTSCore (heading), and LLM as a judge decision.
Ground Truth Gen. Lim. GPT 4 Llama 3 User
**Model Complexity Concerns**: - There is a question regarding whether the performance gains are due to increased model complexity rather than the proposed recursive mixing approach. **Potential for Increased Complexity**: Although the method is described as simple, the recursive nature of the approach may introduce complexity in implementation and understanding, particularly for practitioners who may not be familiar with the underlying concepts Yes Yes Yes
**Computational Complexity**: The Dual-aspect Attention mechanism significantly increases computational complexity, which may affect the performance and practicality of the proposed method. ’**Complexity of the Model**: The DACT arch, while innovative, introduces additional complexity compared to traditional models’, ’This complexity may lead to longer training times and increased computational resource requirements, which could be a barrier for practical applications in resource-constrained environments’ Yes Yes Yes
**Insufficient Detailed Experimental Analysis**: The paper primarily provides quantitative results without a detailed experimental analysis, which limits the understanding of the findings. **Limited Discussion on Failure Cases**: While the paper discusses success and failure cases of existing methods, it may not provide sufficient insight into the specific conditions under which SCILL might fail, limiting the understanding of its robustness. Yes No Yes
**Connection to Practical Methods**: There is an unknown connection between the theoretical findings and popular ensemble methods used in practice, which raises questions about the practical relevance of the work. **Unexplored Variants of Ensemble Methods**: The research does not explore other ensemble methods beyond those mentioned, such as boosting or stacking, which may have different theoretical properties and could provide additional insights into the interpolating regime No Yes Yes
**Lack of Exploration of Alternative Methods**: The paper does not sufficiently explore or compare the performance of other methods like MFVI and SGLD in the context of covariate shift, which could provide a more comprehensive understanding of the problem. **Absence of Dataset Details**: The dataset section is marked as "nan," indicating a lack of information about the datasets used for experimentation’, ’This omission makes it difficult to assess the validity and applicability of the findings, as the choice of dataset can significantly influence the results No No No
Table 23: LLM as a Judge (GPT 4, Llama 3) and User to decide whether Ground Truth Limitation and LLM Generated Limitation have the same topic or not.
Note: Here GPT 4, Llama 3, and User decision regarding both limitation match or not.
Ground Truth Limitations: Our goal is to provide comprehensive study of learning algorithms, including their hyperparameters. However, our study is limited on the variety of other design choices to restrict the experimental complexity. While we studied both synthetic and realistic image datasets, both these datasets are relatively simple with the same small number of generative factors and each of the factor follows uniform distribution. For learning algorithms, we focus on studying the inductive bias on the representation format while fixing the model arch design which can impact the results. Moreover, we did not study hyperparameters beyond those related to the latent representations. Specifically, we did not study how the type and configurations of the optimizer and the batch size would change the results; instead, we followed common setups in previous studies. '''   LLM Refined Limitations: ’1. Here are the limitations extracted from the provided limitation:’, ’2. . **Experimental Complexity**: The study is limited in the variety of design choices to reduce experimental complexity.’, ’3. . **Dataset Complexity**: The study utilizes relatively simple synthetic and realistic image datasets, which have a small number of generative factors.’, ’4. . **Generative Factors**: Each generative factor in the datasets follows a uniform distribution, limiting the diversity of the data.’, ’5. . **Model arch**: The focus is on studying the inductive bias related to the representation format while keeping the model arch design fixed, which may influence the results.’, ’ 6. . **Hyperparameter Scope**: The study does not explore hyperparameters beyond those related to latent representations.’, ’7. . **Optimizer and Batch Size**: The impact of different types and configurations of optimizers and batch sizes on the results is not examined, as the study adheres to common setups from previous research.’
Figure 5: Ground Truth Limitations and LLM Extracted Limitations in NeurIPS dataset.
Ground Truth Limitations: While our work provides useful starting point for understanding student feedback,there are limitations to our work. Addressing these limitations will be an important area for future research. Comments may not reflect real student feedback. The comments in our dataset are from users who have chosen to post publicly on youtube. Addi- tionally,the comments may include features specific to this online education setting. Thus,the comments may reflect real student comments from these courses. There is selection bias in lecture sources. Sight includes lectures that may be drawn from the most successful offerings of that course. The instructional quality may not be representative of typical instruction. Thus,inferences drawn about the instruction should be interpreted with caution,as they might not generalize to other lecture settings. We analyze only english comments. We analyze only english comments because the lecture content is given in english and the authors are most comfortable with english. As result,our rubric may not capture the types of feedback from nonenglish students watching lectures taught in english. we annotate small subsample of the data to assess the validity of the automatic labels,we conduct diagnostic study on small,randomly selected subset of the dataset,comprising approximately of the comments. Our work aims to establish preliminary evaluation of the humanmodel agreement and model annotations,and further validation of the automatic labels is necessary. '''   LLM Refined Limitations: Here are the limitations extracted from the provided limitation: 1. **Comments May Not Reflect Real Student Feedback**: The dataset consists of comments from users who have chosen to post publicly on YouTube, which may not accurately represent genuine student feedback. 2. **Features Specific to Online Education Setting**: The comments may include characteristics that are unique to the online education context, potentially skewing the interpretation of student feedback. 3. **Selection Bias in Lecture Sources**: The lectures included in the analysis may be drawn from the most successful offerings of a course, leading to a selection bias that does not represent typical instructional quality. 4. **Generalizability of Instructional Quality**: Inferences about instructional quality based on the selected lectures should be made with caution, as they may not generalize to other lecture settings. 5. **Analysis Limited to English Comments**: The analysis is restricted to comments in English, which may overlook feedback from non-English speaking students who are watching lectures taught in English. 6. **Small Subsample for Validity Assessment**: The study annotates a small subsample of the data to assess the validity of automatic labels, which may not provide a comprehensive evaluation of the entire dataset. 7. **Need for Further Validation of Automatic Labels**: The work establishes a preliminary evaluation of human-model agreement and model annotations, indicating that further validation of the automatic labels is necessary.
Figure 6: Ground Truth Limitations and LLM Extracted Limitations in ACL dataset.
Ground Truth OpenReview: Reviewer 1: This paper tries to tackle the domain generalization problem by finding flat minima. It is interesting to investigate the loss landscape of the DNNs on domain generalization problems. Originality: The method proposed in this paper is mainly built on previous approaches such as SWA, and the authors mention this in the submission. Quality: This submission is complete piece of work and the claims are well supported by the empirical results. The evaluation is reasonable. Clarity: This submission is well-written and easy to follow. Significance: This paper provides interesting empirical observations and improves the state-of-the-art by applying their proposed new approach. Pros: The proposed method consistently improves the model performance on domain generalization problems, and the proposed method is flexible and can be easily integrated with other training approaches for better performance. Cons: The generalization bound does not provide intuition for robust risk minimization.. Reviewer 2: Originality The paper is, perhaps by design, not particularly original. Rather, it makes the observation that an existing set of techniques with particular goal also has applications to different goal. This is still, to the best of my knowledge, novel observation and contribution, and furthermore the modifications made to SWA to arrive at SWAD also appear novel. Nevertheless, these modifications are relatively small, so the method can still be viewed as version of the general idea of SWA, and thus originality is not the paper’ strong suit. Quality The paper is of relatively high quality, primarily due to the empirical results. did not carefully check the theory for correctness, but am unsure as to the significance of the theorems. They appear rather similar to prior theoretical results on how flat minima can lead to better generalization, but with some additional divergence terms thrown in as we are dealing with the DG setting… Reviewer 3: Pros: The paper tackles an important and relevant problem to the ML community. The proposed algorithm is relatively simple to implement in practice and perform well experimentally. Cons: My major concern is wrt to the novelty of the work and the connection/motivation between the proposed algorithm SWAD, domain generalization and the derived theoretical insights… '''   LLM Refined OpenReview: "1. . **Connection Between Theory and Method**: The relationship between the theoretical framework presented in the paper and the SWAD algorithm is weak. There is a need for a clearer connection between the flatness term in the equations and the algorithm’s performance.", ’2. . **Novelty of Theoretical Results**: The novelty of the theorem compared to classic results, such as those from Ben-David, is questioned. The theoretical contributions do not significantly advance the understanding of flat minima in the context of domain generalization.’, ’3. . **Generalization Bound Intuition**: The generalization bound provided does not offer sufficient intuition for robust risk minimization (RRM), and it does not convincingly demonstrate that RRM is superior to standard empirical risk minimization (ERM).’], ’4. . **Dependence on Previous Work**: The proposed method heavily relies on the existing SWA approach, with only minor modifications. This raises concerns about the originality of the contribution.’, ’5. . **Theoretical Framework Limitations**: The reliance on VC dimension in the theoretical framework is problematic, as it can be excessively large for the models used, especially in domain generalization scenarios.’, ’6. . **Overlooked Details in Proofs**: There are concerns that important terms have been overlooked in the technical proofs, which could undermine the validity of the results.’, ’7. . **Lack of Clear Motivation for SWAD**: The motivation for using SWAD over SWA is not clearly articulated, making it difficult to understand the advantages of the proposed algorithm.’, ’8. . **Introduction of New Hyperparameters**: The proposed algorithm introduces additional hyperparameters, which may complicate the implementation and tuning process.’, ’9. . **Limited Societal Impact Discussion**: While a limitations section is included, there is a lack of discussion regarding the societal impacts of the research, although specific concerns were not identified.’, "10. . **Potential for Broader Applications**: The method’s applicability beyond domain generalization is not explored, which could limit its significance and impact in the broader machine learning community.
Figure 7: Ground Truth and LLM Extracted Limitations in OpenReview.

### v2

@@
 Abstract
-In scientific research, “limitations” refer to the shortcomings, constraints, or weaknesses within a study.
+In scientific research, “limitations” refer to the shortcomings, constraints, or weaknesses of a study.
 A transparent reporting of such limitations can enhance the quality and reproducibility of research and improve public trust in science.
-However, authors often a) underreport them in the papers’ text and b) use hedging strategies to satisfy editorial requirements at the cost of readers’ clarity and confidence.
-This underreporting behavior, along with an explosion in the number of publications, has created a pressing need to automatically extract/generate such limitations from scholarly papers.
-In that direction, we report a complete architecture for computational analysis of research limitations.
-Specifically, we a) create a dataset of limitations in ACL, NeurIPS, and PeerJ papers by extracting them from papers’ text and integrating them with external reviews; b) propose methods to automatically generate them using a novel Retrieval Augmented Generation (RAG) technique; c) create a fine-grained evaluation framework for generated limitations and provide a meta-evaluation for the proposed evaluation techniques.
-Code and datasets are here: github | huggingface
-[…]
+However, authors often underreport limitations in their papers and rely on hedging strategies to meet editorial requirements at the expense of readers’ clarity and confidence.
+This tendency, combined with the surge in scientific publications, has created a pressing need for automated approaches to extract and generate limitations from scholarly papers.
+To address this need, we present a full architecture for computational analysis of research limitations.
+Specifically, we (1) create a dataset of limitations from ACL, NeurIPS, and PeerJ papers by extracting them from the text and supplementing them with external reviews; (2) we propose methods to automatically generate limitations using a novel Retrieval Augmented Generation (RAG) technique; (3) we design a fine-grained evaluation framework for generated limitations, along with a meta-evaluation of these techniques.
+Code and datasets are available at: Code: https://github.com/IbrahimAlAzhar/BAGELS_Limitation_Gen Dataset: https://huggingface.co/datasets/IbrahimAlAzhar/limitation-generation-dataset-bagels
+Accepted to the Findings of the 2025 Conference on Empirical Methods in Natural Language Processing (EMNLP 2025).
 BAGELS: Benchmarking the Automated Generation and Extraction of Limitations from Scholarly Text
-Ibrahim Al Azher, Miftahul Jannat Mokarrama, Zhishuai Guo, Sagnik Ray Choudhury, Hamed Alhoori Northern Illinois University, Dekalb, IL, USA University of North Texas, Denton, TX, USA {iazher, mmokarrama1, zguo, alhoori}@niu.edu, sagnik.raychoudhury@unt.edu
+Ibrahim Al Azher, Miftahul Jannat Mokarrama, Zhishuai Guo, Sagnik Ray Choudhury, Hamed Alhoori Northern Illinois University, DeKalb, IL, USA University of North Texas, Denton, TX, USA {iazher1, mmokarrama1, zguo, alhoori}@niu.edu, sagnik.raychoudhury@unt.edu
 1 Introduction
-Figure 1: System Architecture.
-We build a dataset by extracting limitations from a paper’s text (when mentioned) and reviews, build a generator for papers that do not explicitly mention limitations, and develop a novel framework for reference-based evaluation of generated limitations.
-In the context of a scientific article, “limitations” refer to the inherent shortcomings, constraints, or weaknesses within a study that may influence the results or limit the generalizability of its findings Ross and Bibler Zaidi (2019).
-These limitations can arise from various elements of the research process, including the methodology, theoretical framework, data collection, experimentation, and analysis Ioannidis (2007).
-Limitations explicitly acknowledged by the authors often include internal validity issues, measurement errors, potential confounding factors, and the failure to measure important variables Puhan et al.
+Figure 1: System architecture for dataset creation, limitation generation, and evaluation.
+In scientific articles, “limitations” refer to the inherent shortcomings, constraints, or weaknesses of a study that may influence its results or restrict the generalizability of its findings Ross and Bibler Zaidi (2019).
+Such limitations can arise from various aspects of the research process, including the methodology, theoretical framework, data collection, experimentation, and analysis Ioannidis (2007).
+Authors commonly acknowledge issues such as internal validity concerns, measurement errors, confounding factors, and the omission of important variables Puhan et al.
 (2009).
-Openly discussing limitations is crucial for several reasons.
-It ensures the credibility and ethical standards of a scientific inquiry Bunniss and Kelly (2010); Chasan-Taber (2014); Annesley (2010).
-By clearly defining the boundaries of a study, researchers provide a realistic context for their findings, supporting accurate interpretation, transferability, and reproducibility Ioannidis (2007); Eva and Lingard (2008).
-Acknowledging limitations fosters scientific integrity by demonstrating a commitment to honesty Žydžiūnaitė (2018).
-It also opens pathways for refining research methods, opening areas for future work Azher et al.
-(2025), and leading to innovations that help future researchers avoid repeating the same shortcomings Escande et al.
+Openly discussing limitations is crucial.
+It upholds credibility and scientific integrity by demonstrating a commitment to ethical and transparent research practices Bunniss and Kelly (2010); Chasan-Taber (2014); Annesley (2010); Žydžiūnaitė (2018).
+It also clarifies the scope of a study, supporting accurate interpretation, transferability, and reproducibility Ioannidis (2007); Eva and Lingard (2008).
+In addition, it helps researchers avoid repeating the same shortcomings Escande et al.
+(2016) while creating opportunities to refine methods and guide future research Azher et al.
 (2025).
-Despite the benefit of honest admission of such limitations, researchers sometimes remain reluctant to include them in the text or articulate them in detail Ioannidis (2007); Ter Riet et al.
+Despite these benefits, researchers are often reluctant to include limitations or articulate them in detail Ioannidis (2007); Ter Riet et al.
 (2013).
-Concerns about the potential impact on publication chances and, subsequently, career progression Montori et al.
-(2004) can encourage this biased behavior.
-When forced to explicitly acknowledge limitations in the text [^1] As is the practice in the NLP and ML community in recent years., authors sometimes resort to using generic or irrelevant terms, which do not meaningfully inform the reader about the study’s real constraints Ross and Bibler Zaidi (2019).
-Moreover, limitations can sometimes serve as a form of “hedging”, where authors present their findings cautiously to avoid making definitive statements Hyland (1998).
-This approach, while safer for the authors, does not contribute to the clarity or usefulness of the research.
-Failure to disclose limitations undermines the scientific process and misleads readers, reviewers, and policymakers, potentially preventing them from recognizing the constrained nature of findings and possible biases Greener (2018).
-NLP research in automatic extraction, generation, and evaluation of limitations from scholarly papers is limited due to the lack of standardized datasets, novel methods, and evaluation framework.
-This study takes a step toward bridging this gap.
-The number of scientific publications has exploded recently Larsen and Von Ins (2010).
-That, combined with the problems of self-reporting, has necessitated a need for computational methods to study the limitations of scholarly work.
-In that direction, our contributions are as follows (see Figure 1):
-  • We create a dataset for limitations reported in research articles.
-For each paper, we extract the limitations from the papers’ text and OpenReview feedback  [^2] https://openreview.net/.
-This creates a valuable benchmark for future studies that want to extract and analyze limitations from research papers and study their impact on scholarly progress.
-Notably, due to the integration of authored text and external data, this benchmark suffers less from the self-reporting bias, providing a broader and more nuanced perspective about limitations.
-  • We create a novel RAG system to generate limitations. This is a useful tool for editors or researchers in the under-reporting or hedging scenario.
-  • Evaluating the generated limitations is a key challenge.
-Many traditional metrics for measuring text generation quality depend on lexical overlap (ROUGE Lin (2004), BLEU Papineni (2001)) or semantic similarity (BERTScore Zhang et al.
+Concerns about the potential impact on publication chances and career progression Montori et al.
+(2004) can reinforce this tendency.
+Even when required to acknowledge limitations, as is now common in NLP/ML research, authors sometimes resort to generic or irrelevant statements that obscure the study’s real constraints Ross and Bibler Zaidi (2019).
+Moreover, limitations may serve as a form of hedging, where findings are presented cautiously to avoid making definitive claims Hyland (1998).
+This practice, while safer for authors, reduces the clarity and usefulness of the research.
+Failure to disclose limitations undermines the scientific process and misleads readers, reviewers, and policymakers, preventing recognition of constrained findings and potential biases Greener (2018).
+Meanwhile, the volume of scientific publications has surged Bornmann et al.
+(2021).
+These factors highlight the need for computational methods to study research limitations.
+However, progress in NLP toward automatic extraction, generation, and evaluation of limitations remains limited, largely due to the lack of standardized datasets, novel methods, and robust evaluation frameworks.
+This study takes a step toward closing this gap.
+Our contributions are as follows (see Figure 1):
+  • Dataset creation.
+We build a dataset of research limitations by extracting them from papers and their reviews.
+By integrating author-reported and reviewer-identified limitations, this benchmark reduces self-reporting bias and provides a broader, more reliable resource for analyzing limitations and their impact on research.
+  • Limitation generation. We design a novel RAG system to automatically generate limitations, offering a way to supplement papers with high-quality, context-aware limitation statements.
+  • Evaluation framework.
+We introduce a new evaluation paradigm for generated limitations.
+Unlike traditional metrics (e.g., ROUGE Lin (2004), BLEU Papineni (2001), BERTScore Zhang et al.
 (2019), MoverScore Zhao et al.
-(2019)).
-They are not sufficient for our purpose as they overprioritize common terms (e.g., tokens such as “bias”, “dataset”, and “generalizability” are standard in the limitations section of most AI/ML/NLP papers and should not affect the evaluation score).
-We develop a new evaluation framework that uses LLMs-as-judges to provide fine-grained comparisons and detailed error analysis, thus offering a more transparent and actionable evaluation for limitations.
+(2019)), which overemphasize common terms (e.g., bias, dataset, and generalizability), our framework leverages LLMs-as-judges for fine-grained, interpretable assessments and actionable error analysis.
 2 Related Work
-Various studies have studied the reporting, or lack thereof, of ‘limitations’ in scholarly papers.
-Ioannidis (2007) found that only 17% of top-tier journal articles mentioned limitations, with just 1% in abstracts.
+Several studies have examined how limitations are reported in papers.
+Ioannidis (2007) found that only 17% of top-tier articles mentioned limitations, with just 1% doing so in abstracts.
 Similarly, Puhan et al.
-Puhan et al.
-(2012) identified 27% of biomedical papers lacked limitations, risking overestimated research reliability.
-Moreover, Goodman Goodman et al.
-(1994) found that in the peer review process acknowledging limitations is often problematic.
-Few journals require discussing limitations Ioannidis (2007), risking biased reviews and weakening scientific dialogue Horton (2002), necessitating reforms for transparency.
-These gaps underscore the need for a systematic framework for limitations generation, an area that remains underexplored.
+(2012) reported that 27% of biomedical papers lacked limitations, risking overestimation of research reliability.
+Goodman et al.
+(1994) noted that acknowledging limitations is often problematic in peer review.
+Few journals require discussing limitations Ioannidis (2007), which can bias reviews and weaken scientific dialogue Horton (2002), highlighting the need for greater transparency.
+Recent work has explored computational approaches to research limitations.
 Faizullah et al.
-Faizullah et al.
-(2024) propose an LLM-chain pipeline to summarize and iteratively refine candidate limitations; Al Azher et al.
-(2024) integrates topic modeling with LLMs to derive and elaborate structured limitation themes; and Azher (2024) develops a graph-augmented LLM approach for generating comprehensive limitation statements.
-Other work focuses on overcoming limitations of charts and graph by generating meanigful captions Al Azher and Alhoori (2024).
-However, these studies are confined to ACL/EMNLP corpora and author-stated limitations, overlooking reviewer comments (e.g., OpenReview), and evaluate only with ROUGE, BERTScore, and coherence metrics, which do not assess finer-grained topic or contextual alignment.
-Our proposed framework addresses existing limitations by incorporating NeurIPS papers alongside ACL, collecting OpenReview comments for each NeurIPS paper to establish a broader ground truth, and implementing a limitation-level text evaluation method to preserve granularity.
-Evaluating NLP outputs is crucial for assessing quality, accuracy, and relevance.
+(2024) proposed an LLM-chain pipeline to summarize and refine candidate limitations.
+Al Azher et al.
+(2024) integrated topic modeling with LLMs to derive structured limitation themes.
+Al Azher (2024) developed a graph-augmented LLM method for generating detailed limitation statements.
+Other studies address the shortcomings of visualizations by generating more meaningful captions for charts and graphs Al Azher and Alhoori (2024).
+However, these studies are limited to ACL/EMNLP corpora and rely on author‐stated limitations, and use metrics such as ROUGE and BERTScore that miss finer-grained contextual alignment.
+Concurrent to our work, Xu et al.
+(2025) also address these gaps by introducing LIMITGEN, a benchmark that incorporates human-written peer reviews to systematically evaluate how well LLMs identify limitations.
+Our framework complements this effort by additionally leveraging cited papers for broader context and introducing a novel limitation-level evaluation method to preserve granularity.
+Evaluating NLP outputs is essential for assessing quality, accuracy, and relevance.
 Traditional metrics like ROUGE and BLEU struggle with semantics, while BERTScore improves similarity but relies on references and lacks meaningful error analysis.
-Recent advancements in large language models (LLMs) have revolutionized text evaluation methods Zheng et al.
-(2023) using zero-shot and in-context learning Wei et al.
+Advances in large language models (LLMs) have opened new evaluation avenues Zheng et al.
+(2023), from zero-shot and in-context learning Wei et al.
+(2022) to specialized approaches such as GPTScore Fu et al.
+(2023), TIGERScore Jiang et al.
+(2023), and PandaLM Wang et al.
 (2023).
-For example, GPTScore Fu et al.
-(2023) uses LLMs to evaluate how likely a text could be generated based on a given context, TIGERScore Jiang et al.
-(2023) provide explainable error analysis, and PandaLM Wang et al.
-(2023) distinguish superior models among multiple models.
-Other approaches, such as AttrScore Yue et al.
-(2023), evaluate whether a reference supports or contradicts generated statements.
-SummacConv Laban et al.
-(2022) breaks documents into sentence-level units and removes summary sentences with low entailment probability.
-Despite their promise, LLM-based evaluations have issues such as positioning bias, where altering the order of inputs affects results.
-To mitigate positioning bias, we systematically alter the position of limitations and select outputs that remain consistent across different orderings.
-Moreover, while these text evaluation methods focus on lexical overlap and surface-level LLM-based analysis, our proposed approach advances beyond these limitations by uniquely combining granularity-aware evaluation, considering LLM as a judge, quantitative scoring, with topic-level agreement.
-3 Limitation Extraction
-Our limitation extraction is twofold.
-At first, we extracted limitations with a science parse  [^3] https://github.com/allenai/science-parse tool.
-If a paper has a separate limitation section, we were able to extract it using this tool.
-But if the limitation has no separate section, we used Python regex to check the word ‘limitation’ and extracted all consecutive texts before another section appeared.
-However, this approach also collected some noisy sentences unrelated to the limitation.
-So, in the later part, we applied LLM and sent the texts to extract only limitations.
-We also used Selenium to collect reviews from OpenReview.
-Later, the ‘author mentioned limitation’ is sent to the zero-shot LLM to extract limitations, and the review from ’OpenReview’ is sent to the zero-shot LLM to extract shortcomings/weaknesses.
-Through this process, we removed the noisy sentences.
-(§ 3.1)
+Other methods include AttrScore Yue et al.
+(2023), which checks factual support, and SummacConv Laban et al.
+(2022), which filters low-entailment sentences.
+Despite their promise, LLM-based evaluations face issues such as positioning bias, where input order can shift results.
+We address this by randomizing order and retaining stable outputs.
+More broadly, our evaluation advances beyond prior work by combining granularity-aware scoring, topic-level agreement, and LLMs-as-judges.
+Taken together, prior research shows both the need and the opportunity for a more systematic treatment of research limitations.
+Building on these insights, our work unifies dataset construction, limitation generation, and evaluation into a single framework, laying the foundation for more transparent and reproducible analysis of limitations.
+3 Limitation Extraction & Evaluation
 3.1 Dataset of Extracted Limitations
 Granularity.
-A key challenge in constructing a dataset of research limitations is defining the appropriate level of granularity.
+A key challenge in building a dataset of research limitations is defining the appropriate level of granularity.
 Should a limitation be captured as a single phrase, a full sentence, or an entire paragraph?
@@
 Although author-reported limitations often provide well-structured insights, previous research indicates that such limitations may be underreported or carefully hedged.
-To address this gap, we incorporate OpenReview [^4] https://openreview.net/ comments, in which peer reviewers frequently identify additional constraints or weaknesses not mentioned by the authors.
-Our dataset consists of papers from conferences in natural language processing and machine learning: ACL [^5] https://aclrollingreview.org/cfp, NeurIPS [^6] https://neurips.cc/public/guides/PaperChecklist, and PeerJ  [^7] https://peerj.com/benefits/indexing-and-impact-factor/.
-We include 6,932 NeurIPS papers from 2021 and 2022 and 5,739 ACL papers from 2023-2024, 1000 biological and medical science papers from PeerJ.
-In addition, we integrate OpenReview comments for 2,802 papers from NeurIPS, thus capturing both self-reported limitations and those externally identified by reviewers.
-All of the PeerJ paper contains self-reported Limitations alongside other sections and peer review comments.
-The first experiment evaluates state-of-the-art LLMs’ ability to extract limitations from papers (when they are explicitly mentioned) and their reviews.
-For each paper, we use GPT-4o mini to extract and get an average of $8$ limitations from a paper and $10$ from open review.
-LLM as a Ground Truth Limitation Extractor.
-Identifying and isolating limitations can be challenging, as they often appear in sections such as Conclusion, Discussion, or Future Work and may also be dispersed throughout OpenReview feedback.
-To address this, we implement a targeted extraction process designed to be both efficient and accurate.
-Rather than processing entire papers, which would be computationally expensive due to larger context windows, we first apply Python regex to identify and extract relevant sections.
-Some research papers include a dedicated limitations section; we extract the text of this section using the AllenAI Science Parse tool [^8] https://github.com/allenai/science-parse.
-When a paper doesn’t have such a well-defined section, we identify sentences containing the word limitations and continue extracting subsequent sentences until we encounter one of the following sections: Conclusion, Future Work, Ethics, Acknowledgement, or Grant.
-Then, we extracted the limitations from the research paper by LLM to remove noisy sentences and make segments.
-To capture broader perspectives from peer reviews, we aggregate comments from multiple OpenReview responses into a single text.
-Finally, we employ GPT-4o  [^9] https://openai.com/index/hello-gpt-4o/ (Figure 4) to segment the extracted text and identify distinct limitation statements, distinguishing those reported by authors from those highlighted by reviewers.
-We evaluate the quality of these extracted limitations through a user study (§5).
-Dataset Applications.
-The resulting dataset will be made publicly available and can be used as a benchmark for evaluating automated limitation extraction and generation methods (§5).
+To address this gap, we incorporate review comments, where reviewers often highlight additional constraints or weaknesses not mentioned by the authors.
+Our dataset includes papers from major NLP and ML conferences, including ACL [^1] https://aclrollingreview.org/cfp and NeurIPS [^2] https://neurips.cc/public/guides/PaperChecklist, as well as biomedical research from PeerJ  [^3] https://peerj.com/benefits/indexing-and-impact-factor/.
+We collect 6,932 NeurIPS papers (2021-2022), 5,739 ACL papers (2023-2024), and 1000 papers from PeerJ.
+In addition, we integrate OpenReview  [^4] https://openreview.net/ comments for 2,802 papers from NeurIPS.
+All of the PeerJ papers contain self-reported limitations alongside other sections and peer review comments.
+For each paper, we use LLM to extract and get an average of $8$ limitations from a paper and $10$ from their reviews.
+3.2 Extraction Process
+We extract spans (blocks) of text from papers or review comments, and then refine them with LLMs, as opposed to passing in the entire paper to an LLM.
+This strikes a balance between accuracy and LLM usage cost.
+1. Limitation Span Extraction: This step extracts blocks of text from the papers that correspond to limitations. We consider both explicit and implicit limitation statements:
+a.
+Explicit limitations.
+These appear in a dedicated limitations section or subsection.
+We identify them using the AllenAI Science Parse tool  [^5] https://github.com/allenai/science-parse, which segments papers into a structured JSON format, allowing for direct and reliable extraction of these dedicated sections.
+For peer review content in NeurIPS papers, we used Selenium to scrape the main review field from OpenReview, which typically includes both strengths and weaknesses of a paper.
+b.
+Implicit limitations.
+These are embedded in broader sections such as discussion or conclusion.
+To identify them, we apply a Python regex script that searches for keywords such as limitation(s), or shortcoming(s).
+To improve precision, we exclude sections where limitations are rarely discussed (e.g., abstract, introduction, related work).
+Our script begins extraction when a limitation-related keyword is detected and continues until a terminal section marker is reached; extraction stops at terms such as acknowledgements, grant, future work, discussion, conclusion, or appendix.
+Although this process is effective, the regex approach for implicit limitations can occasionally capture irrelevant sentences, introducing noise into the results.
+2.
+Refinement via LLM: To improve precision, we use an LLM to filter meaningful limitations from the tool-extracted ones (from both papers and review) by removing noisy sentences.
+Importantly, we strictly instruct the LLMs to extract limitation statements without paraphrasing, altering, or generating new content, and producing them as a structured sequence of sentences, denoted by $L_{i}=\{l_{i1},l_{i2},\dots,l_{ix}\}$.
+To incorporate broader perspectives from peer reviews, we first aggregate comments from multiple review responses into a single consolidated text.
+We then prompt the LLMs (Figure 3, appendix) to segment this text and identify distinct limitation statements by reviewers, with the latter being denoted as $R_{i}=\{r_{i1},r_{i2},\dots,r_{ix}\}$.
+Following this extraction, a master LLM is tasked with merging the author-reported limitations $L_{i}$ and the reviewer-identified limitations $R_{i}$ of input paper $P_{i}$.
+The model is explicitly instructed to merge only those limitation statements that were identical or semantically equivalent across both the author-mentioned limitations and the peer review.
+As before, the model is restricted from changing, rephrasing, or reordering any sentences during the merge process, and we get final Ground truth extracted limitations $G_{i}=\{g_{i1},g_{i2},\dots,g_{ix}\}$.
+Finally, we evaluate the quality of these extracted and merged limitations through a user study described in § 3.3.
+We use GPT 4o-mini as both the extractor and master LLM (Examples of limitations extraction by LLM from NeurIPS, ACL, and OpenReview are provided in the Appendix in Figure 6, 7, and 8, respectively).
+3.3 Limitation Extraction Evaluation
+Are the limitations extracted or generated?
+The first goal in the evaluation process is to check if the LLM extracted limitations are grounded in the text, i.e., they only come from the input (papers/reviews) and not from the LLMs’ parametric knowledge or hallucinations.
+For this, we employ three annotators (separate from this paper’s authors) [^6] CS graduate students with research experience in NLP and AI.
+The first ground truth consists of only author-mentioned limitations.
+We choose a sample of 100 limitations from ACL, NeurIPS, and PeerJ, and for each, we show them the source and ask a Yes/No question, whether they thought the LLM extracted the limitation from the source without generating text.
+Each annotator answer positively in $90\%$ of cases (avg $\pm$ std= $95\pm 2.45\%$) (Table 1).
+Model Role Sample U1 U2 U3
+GPT 4o-mini Extractor 100 92 95 98
+Table 1: Evaluating LLM as an extractor role with human annotator (U).
+In the second evaluation, two annotators manually verified the extracted limitations from 1000 papers from NeurIPS and PeerJ and their reviews.
+The annotators assessed whether 1) each LLM-extracted author mentioned limitation was grounded in the source paper, (2) each extracted limitation from the peer review was also grounded in the review, and (3) the merged set (limitation + review) included only truly overlapping or matching limitations between the two sources.
+Their analysis confirmed that all extracted limitations were faithfully sourced, with no instances of hallucinated, noisy, or newly generated content.
+We also computed the performance of the Llama3 70B for this extraction task, and the result was unsatisfactory.
+The quality of the extraction.
+The SMEs from the last step annotated 500 ACL papers and 100 NeurIPS papers: one annotator extracted limitations (taking the full section when explicit, or selecting limitation-related sentences when implicit), and two others verified the results.
+We then compared the tool-based (GPT-4o mini) extractions against this gold standard.
+Notably, the human-extracted (gold) limitations were not segmented; therefore, we combined the LLM-extracted limitations and compared them with the gold ones using cosine similarity, precision, recall, F1, and fuzzy matching [^7] These strings are tokenized.
+(Table 2): ACL achieved a strong F1 of 85.69, likely aided by more frequent explicit limitation sections.
+NeurIPS yielded a moderate F1 of 72.42, reflecting the more scattered, implicit presentation of limitations where LLM should be utilized to remove noisy information.
+Dataset CS P R F1 Fuzzy
+ACL 89.38 89.63 84.93 85.69 91.18
+Neurips 78.08 68.76 84.13 72.42 70.26
+Table 2: Performance between Human Extracted Limitations vs Tool Extracted Limitations in Cosine Similarity (CS), Precision (P), Recall (R), F1 score (F1), and Fuzzy matching
+3.4 Dataset Applications
+The resulting dataset is publicly available [^8] https://huggingface.co/datasets/IbrahimAlAzhar/limitation-generation-dataset-bagels and can be used as a benchmark for evaluating automated limitation extraction and generation methods (§5).
 Beyond this, the extracted limitations can be examined and organized into a taxonomy of limitations in ML and NLP, offering a more structured understanding of common research challenges.
 By integrating this taxonomy into citation networks, we can introduce the concept of a Limitation Multigraph, enabling scientometric analyses into whether certain limitations shape the direction of subsequent research or, alternatively, tend to be overlooked.
-These avenues present new opportunities to study how the reporting (or lack thereof) of limitations affects the broader scientific discourse, a topic we plan to explore in future work.
-Figure 2: Evaluation of generated limitations.
+These avenues present new opportunities to study how the reporting (or the lack thereof) of limitations affects the broader scientific discourse, a topic we plan to explore in future work.
 4 Limitation Generation
-We generated limitations by feeding every paper’s text, excluding its actual “Limitations” section into both vanilla and RAG-augmented LLMs, experimenting with either the top three sections (Abstract, Introduction, Conclusion) or the full paper as input and using three ground-truth definitions (author-stated limitations, OpenReview comments, or their combination).
-We computed each section’s cosine similarity to the “Limitations” section across ACL, NeurIPS dataset and found that Abstract, Introduction, and Conclusion rank highest on average (Table 19, Appendix).
-When full paper inputs exceed the context length, we make them chunkwise.
-4.1 LLM and RAG for Limitation Generation
-Most research papers do not explicitly mention limitations, or they underreport these shortcomings even when a dedicated section is provided.
-To address this gap, we propose a system that generates limitations directly from the paper text.
+Most research papers either do not explicitly mention limitations or underreport them, even when a dedicated section is provided.
+We compare two systems’ ability to generate limitations from research papers: (a) vanilla LLM and (b) RAG.
+Note that the generators don’t have access to the text from where the limitations are extracted, e.g., limitation sections of the papers, paragraphs identified as limitations, or paper reviews; otherwise, the task would be trivial.
+To improve computational efficiency, we use the three most important sections of a paper as input to the generators rather than the full text.
+The importance score is computed by the cosine similarity of a section and a reference limitation embedding (see Table 8, Appendix).
 Vanilla LLM.
-In the vanilla LLM setup, the language model is prompted with the entire paper $P_{i}$.
-When the paper exceeds the context window, it is divided into chunks $\{P_{i}^{\prime}\}$, and limitations are generated for each chunk D’Arcy et al.
+In the vanilla LLM setup, when the input exceeds the context window, it is divided into chunks $\{P_{i}^{\prime}\}$, and limitations are generated for each chunk D’Arcy et al.
 (2024).
-The LLM then aggregates these chunk-specific outputs into a cohesive, meaningful final set of limitations, ensuring that no relevant material is overlooked in especially long or detailed papers.
+The LLM is then also asked to aggregate these chunk-specific outputs into a cohesive, meaningful final set of limitations.
 RAG Integration.
@@
 In particular, even when a paper lacks an explicit limitations section, other papers with similar methodologies or datasets may discuss relevant shortcomings.
-More importantly, certain findings may be implicitly contradicted by subsequent research.
-Also, failing to consult the broader literature leads to a significant gap in the reported limitations.
-To address this issue, we employ a RAG framework, which allows the system to draw on context from multiple papers rather than relying only on $P_{i}$.
-Specifically, we begin by splitting 100 random papers into multiple chunks and inserting these chunks into a vector database.
-A retriever then identifies the top $k$ chunks most relevant to the input text $\{P_{i}/P_{i}^{\prime}\}$, depending on whether the paper $P_{i}$ or a chunk $P^{\prime}_{i}$ can fit into the context window, using cosine similarity to match these chunks to the query (Figure 1, step 4,5).
+For example, a paper can use SVM and not explicitly mention the modeling assumptions, whereas a related paper possibly will.
+Moreover, certain findings may be implicitly contradicted by subsequent research.
+To address this issue, we employ a RAG framework, which allows the system to draw context from multiple papers rather than relying only on $P_{i}$.
+There can be multiple notions of relatedness; we compare between two: a) relatedness induced by the citation network of $P_{i}$, and b) textual similarity between $P_{i}$ and other papers.
+For the citation network, we use both the Pcited-by, i: papers citing $P_{i}$, and Pcited-in, i: papers that $P_{i}$ cites.
+We parse the reference section of $P_{i}$ to extract the DOI and title of each “cited in” work.
+The “cited by” DOI and titles are collected from the OpenAlex API  [^9] https://openalex.org/.
+We query the Semantic Scholar API  [^10] https://www.semanticscholar.org/product/api with $P_{i}$’s title to get the DOIs for top 5 most semantically close papers.
+These DOIs and titles are cross-referenced with arXiv metadata, and the full texts of the matched papers are downloaded and parsed with the Science Parse tool.
+For each paper $P_{i}$, we build separate RAG indices with a) Pcited-by, i b) Pcited-in, i, c) semantically close papers, and their combinations, where papers are split into chunks by section to preserve detail.
+We combine the strengths of both keyword-based (BM-25) and semantic (FAISS) search by assigning a 50% weight to the scores from each retriever.
+We use a LLM-based reranker, where we retrieve 20 chunks, and pass these chunks to a GPT 4o-mini model along with the original input paper.
+The model is prompted to score the relevance of each chunk on a scale of 1 to 10.
+Only the chunks that receive a relevance score of 8 or higher are ultimately selected.
+We compare this method with the simple baseline of just using the retrieved chunks in §7.4.
+Figure 2: Evaluation of generated limitations.
 5 Evaluation of Generated Limitations
-We introduce a framework, called PointWise (PW), for evaluating the performance of LLMs in generating limitations for scholarly papers (Figure 2).
+We want to evaluate the quality of the generated limitations by comparing them with the extracted ones.
+Functionally, both the ground-truth and the predictions are a set of text segments.
+NLP metrics like BERTScore, ROUGE, and cosine similarity can yield surface overlaps, providing high scores even when the generated limitations are not appropriate, too generic, or imprecise.
+A possible alternative is to use a holistic LLM-as-Judge approach, where the generated/ground-truth limitations are merged into single text blocks and then compared.
+This lacks the point-level granularity needed for fine-grained analysis.
+We address both these problems by introducing the PointWise (PW) evaluation framework (Figure 2).
 Problem Setup.
@@
 To quantify coverage, we first create all possible pairs of limitations between the sets $G_{i}$ and $H_{i}$. Let
-[…]
-Hence, $|S_{i}|=x\times y$. We then use an LLM as a judge Zheng et al. (2023) to decide if a ground truth limitation $g_{ik}$ and a generated limitation $h_{il}$ are similar in content or topic:
+$S_{i}=\{(g_{ik},\,h_{il})\mid 1\leq k\leq x,\,1\leq l\leq y\}.$
+Hence, $|S_{i}|=x\times y$.
+We then use an LLM as a judge Zheng et al.
+(2023) to decide if a ground truth limitation $g_{ik}$ and a generated limitation $h_{il}$ are similar in content or topic (Figure 5, Appendix):
 $J(g_{ik},h_{il})\;=\;\begin{cases}1,&\text{if }g_{ik}\text{ and }h_{il}\text{ are similar},\\ 0,&\text{otherwise}.\end{cases}$
@@
 $A_{Hi}\;=\;\frac{1}{y}\sum_{l=1}^{y}C_{Hi}(h_{il}).$
-We aggregate these coverage values across all papers by taking their means:
+We aggregate these coverage values across all pa- pers by taking their means:
 $A_{G}\;=\;\frac{1}{n}\sum_{i=1}^{n}A_{Gi},\quad A_{H}\;=\;\frac{1}{n}\sum_{i=1}^{n}A_{Hi}.$
+Finally, we display $A_{G}$ and $A_{H}$ as percentage.
 D. Precision, Recall, and F1.
 We also compute overall precision, recall, and F1 scores. For each paper $P_{i}$:
-$\text{TP}_{i}$ $=|M_{i}|,\quad\text{FP}_{i}$ $=x-\sum_{k=1}^{x}C_{Gi}(g_{ik}),$
+$\text{TP}_{i}$ $=|M_{i}|,$
+$\text{FP}_{i}$ $=x-\sum_{k=1}^{x}C_{Gi}(g_{ik}),$
 $\text{FN}_{i}$ $=y-\sum_{l=1}^{y}C_{Hi}(h_{il}).$
 Here, $\text{TP}_{i}$ (true positives) is the total number of matched pairs; $\text{FP}_{i}$ (false positives) is the number of ground truth limitations not matched by any LLM-generated limitation; $\text{FN}_{i}$ (false negatives) is the number of LLM-generated limitations unmatched by any ground truth limitation.
-True negatives ($\text{TN}_{i}$) are not defined since no “negative” class of limitations is specified.
-If there is one ground truth limitation $g_{i1}$ matches with multiple LLM-generated limitations (and vice versa), True Positive (TP) counts as one.
+True negative $(TN_{i})$ is not applicable in this case, as we do not have a defined negative class.
+If there is one ground truth limitation $g_{ik}$ that matches with multiple LLM-generated limitations (and vice versa), True Positive (TP) counts as one.
 (Details in Appendix A.1)
 5.2 Performance
-After finding matched pairs $(g_{ik},h_{il})\in M_{i}$, we evaluate their quality using text based evaluation (Rouge-1, Rouge-2, BERTSCore, Cosine Similarity), keyword based evaluation, and heading based evaluation (Details in Appendix A.2).
-6 Experimental Setup
-We fine-tuned three sequence-to-sequence models, T5 (512-token window) Raffel et al.
+After identifying matched pairs $(g_{ik},h_{il})\in M_{i}$, we score each pair’s quality using (i) text-based metrics: ROUGE-L, BERTScore, and cosine similarity, and (ii) keyword overlap (Jaccard Similarity).
+Finally, the per-pair scores are averaged.
+Unmatched items are excluded, as our goal is to quantify similarity within aligned pairs rather than coverage (details in Appendix A.2).
+6 Experimental Setup for Generation
+We use three LLMs (GPT-3.5, GPT-4o-mini, and Llama 3.1 8B  [^11] https://huggingface.co/meta-llama/Llama-3.1-8B) in a zero-shot setup for both the vanilla generation and RAG.
+The GPT models are accessed through APIs, and the LLama models are locally deployed with Ollama.
+For the vanilla generation, we also fine-tune three sequence-to-sequence models, T5 (512-token window) Raffel et al.
 (2020), BART (1024 tokens), and Pegasus (1024 tokens) Zhang et al.
 (2020) on a 70 / 30 train–test split.
-All were trained for 3 epochs with a learning rate of $5\times 10^{-5}$, weight decay 0.01, 300 warmup steps, and batch sizes of 4 (train) and 8 (eval), with early stopping; inputs longer than 512 tokens were truncated.
-For zero-shot evaluation, we used, GPT-3.5 (16 K) and GPT-4o (128 K) at temperature 0.1, and Llama 3 8B and Mistral 7B (128 K) via Ollama.
-To incorporate retrieval (RAG), we built a VectorStoreIndex with llama-index  [^10] https://www.llamaindex.ai/., and OpenAI text-embedding-ada-002 embedding model for encoding the source and query documents.
-We took 100 random papers from the dataset and stored them in a vector database.
-These documents were chunked into 800-token segments with 20-token overlaps to preserve context, and the LLM temperature was set to 0.
+All models were trained for 3 epochs with a learning rate of $5\times 10^{-5}$, weight decay 0.01, 300 warmup steps, and batch sizes of 4 (train) and 8 (eval), with early stopping; inputs longer than 512 tokens were truncated.
+For RAG, the vector database is built with llama-index  [^12] https://www.llamaindex.ai/, and the OpenAI text-embedding-ada-002 embedding model for encoding the source and query documents.
 7 Experiments and Results
-7.1 Limitation Extraction Evaluation
-We manually collected limitation sections from 500 ACL papers randomly (human curated).
-We measure the n-gram overlap and semantic similarity between human-curated limitations vs.
-automatically extracted ground truth limitations to evaluate our automatic limitation extraction process.
-As shown in Table 1, our automatically extracted limitation text achieves strong alignment with human-curated ground truth limitations on the ACL ’23 and ACL ’24 datasets, with ROUGE-1, L scores exceeding 85% and BERTScore above 95%.
-Metrics ACL 24 ACL 23
-ROUGE-1 86.68 85.07
-ROUGE-2 83.13 81.68
-ROUGE-L 86.64 85.04
-BLEU 77.70 75.15
-Jaccard Similarity 79.36 77.08
-JS Divergence 23.58 24.88
-Levenshtein 92.54 90.32
-Cosine Similarity 90.05 88.70
-BERTScore (F1) 95.70 95.66
-Table 1: Evaluating the Alignment of Human-Collected vs. Automated Limitation Extraction.
-Human Evaluation.
-We employ three annotators (separate from this paper’s authors) [^11] CS graduate students with research experience in NLP and AI to evaluate the quality of the extracted limitations.
-We choose a sample of 100 limitations, and for each, we show them the source and ask a Yes/No question, whether they thought the LLM extracted the limitation from the source, as opposed to generating it from scratch.
-Each annotator answer positively in $>90\%$ of cases (avg $\pm$ std= $95\pm 2.45\%$) (Table 2).
-The annotator pairs have weighted Kappa scores between $38-75\%$, denoting fair to substantial agreement (Table 18, Appendix).
-We checked Llama 3 70b as an extractor, but the extraction quality was not good.
-Overall, the extraction quality is high, and the annotators agree; therefore, we use the extracted limitations as ground truth in the subsequent evaluations.
-Furthermore, two annotators manually verified the extracted limitations from 1000 papers.
-They assessed whether each automatically extracted limitation (via tool + LLM) was grounded in the original paper text or OpenReview content and whether it included any irrelevant or unrelated information.
-Their analysis confirmed that all extracted limitations were faithfully sourced, with no instances of hallucinated, noisy, or newly generated content.
-Model Role Sample U1 U2 U3
-GPT 4o mini Extractor 100 92 95 98
-Table 2: Evaluating LLM as an extractor role with human annotator (U).
-Model R-L BS JStext CStext CGT CLLM Prec. Recall F1
-T5 19.92 87.81 10.82 31.79 35.48 29.59 0.29 0.31 0.30
-BART 19.43 87.67 10.68 31.91 33.71 30.10 0.30 0.31 0.31
-Pegasus 20.15 87.66 10.71 33.39 29.28 25.27 0.25 0.26 0.26
-Llama 3 25.66 88.30 14.69 40.4 61.38 39.04 0.39 0.50 0.44
-GPT 3.5 24.24 87.08 14.65 43.12 76.62 46.65 0.47 0.67 0.55
-GPT 3.5 + RAG 30.21 90.88 19.47 45.37 39.99 44.66 0.42 0.40 0.41
-GPT 4o-mini 16.57 86.02 8.70 32.29 57.65 19.76 0.20 0.31 0.24
-GPT 4o mini + RAG 23.17 87.33 12.99 39.29 67.13 45.67 0.57 0.45 0.51
-Table 3: Results of models in “Coverage” (Coverage of Ground Truth Limitation (CGT), LLM Generated Limitation (CLLM), Precision, Recall, and F1-score) and “performance” metrics – Rouge-x, BLEU, BertScore (BS), Jaccard (JS) and Cosine (CS) similarity over text (ACL data).
+Evaluation of LLM as Aligner.
+The PointWise evaluation protocol above uses an LLM to determine whether a generated limitation matches or aligns with a ground truth one.
+We evaluate GPT 4o-mini’s reliability in this task.
+A set of 100 positive (as per the model prediction) and 100 negative instances is annotated independently by three human evaluators.
+The human annotators have a Cohen’s $\kappa$ score of >=>= 95% (Table 7, Appendix), which shows that the task is largely unambiguous.
+Cohen’s $\kappa$ between human annotators and model (GPT 4o mini) prediction is 90-95%, showing exceptional agreement.
+In comparison, Llama-3.1 400B shows poor agreement with the human judges (76%-81%), so in subsequent evaluations, we use GPT-4o mini as the aligner.
+See Table 9 in the appendix, for an example alignment.
 7.1 Limitation Generation Evaluation
-Coverage Evaluation.
-Table 3 reports results on ACL data using only author-stated limitations as ground truth.
-We couldn’t collect the OpenReview of ACL papers, so we used the author-mentioned limitations as ground truth.
-GPT-3.5 achieves the highest Coverage of Ground Truth (CGT=76.62) and Coverage of LLM-Generated Limitations (CLLM=46.65).
-Vanilla LLMs tend to produce overly verbose limitations.
-Integrating RAG yields more focused, concise outputs that cover a broader range of topics and align better with ground truth—for both GPT-3.5 and GPT-4o mini.
-In particular, incorporating RAG increases CLLM by 25.91 points over GPT-4o mini.
-For NeurIPS data (Table 4), considering RAG increase CGT, decrease CLLM, which gains F1 score (+0.06).
-Compared to ACL, NeurIPS, and Peerj, the ACL dataset archives the highest CGT, and NeurIPS achieves the highest CLLM (Table 11, Appendix).
-Performance Evaluation.
-Table 3 reports our performance metrics using author-extracted limitations as ground truth.
-Compared to all models, GPT 3.5 + RAG shows the best results.
-Augmenting GPT-3.5 with RAG boosts every measure, ROUGE-L (+5.97) and BERTScore (+3.8).
-GPT-4o mini exhibits the same upward trend when paired with RAG.
-For NeurIPS data (Table 4), considering RAG increased the performance of all metrics except Cosine Similarity.
-Among our datasets, NeurIPS achieves the highest ROUGE, BERTScore, and BLEU largely due to the inclusion of OpenReview feedback (Table 11, Appendix).
-Moreover, our proposed PointWise approach achieves better BERTScore than the traditional way in every model (Table 8, 17, Appendix).
-Metric GPT 4o mini GPT 4o mini + RAG
-Rouge-L 17.07 18.10 (+1.03)
-BERTScore 86.25 86.45 (+0.2)
-JStext 9.12 9.72 (+0.6)
-CStext 35.42 34.71 (-0.71)
-CGT 45.69 61.76 (+16.07)
-CLLM 76.27 59.18 (-17.09)
-F1 score 0.54 0.60 (+0.06)
-Table 4: Coverage and Performance measure using GPT 4o mini + RAG in NeurIPS data. Ground truth is Limitation + OpenReview, and input is the top 3 sections alongside the abstract.
-Evaluation ‘LLM as a Judge’.
-We evaluated our Judge LLM model GPT 4o mini regarding PointWise evaluation approach.
-From each pair of ground truth and LLM generated limitation, we asked LLM whether both are same or not (Figure 3).
-We use the same annotators as before and choose a sample of 200 pairs, 100 positive cases where the LLM predicted a match and 100 negatives.
-We ask humans the same question as the model and calculate the weighted Kappa scores.
-As seen in (Table 9, Appendix), the human experts overwhelmingly agree with GPT, with kappa values ranging between $90$-$95\%$.
-The agreement between human annotators is $>92\%$, indicating a perfect agreement.
-We also experimented with a strong baseline of Llama-3.1 400B, which shows a poor agreement with humans.
+The ground truth contains papers that have a) only self-reported limitations and b) limitations coming from both self-reports and reviews.
+7.2 Author-Mentioned Limitation
+We evaluate the model’s ability to generate self-reported limitations on the ACL part of the dataset, as these papers a) have explicit limitation sections, and b) do not have open-access reviews.
+The results are presented in Table 3.
+Vanilla LLMs and fine-tuned models.
+Zero-shot models outperform trained models in almost all metrics, with GPT-3.5 achieving the best results in coverage metrics, and LLama 3 achieving the best in performance metrics.
+Surprisingly, GPT 4o-mini has a significantly worse performance than other zero-shot models.
+However, the performance metrics are based on n-gram overlap and embedding measures (e.g., ROUGE, BLEU, BERTScore, cosine similarity) that primarily capture surface overlap or shallow semantics and can miss factual correctness and completeness.
+Therefore, we prioritize coverage-based metrics, CGT, CLLM, and F1, and report NLP metrics as secondary diagnostics.
+RAG.
+Since GPT-3.5 performs the best in the coverage metrics, we utilize it in a RAG setup, where the index consists of “cited-in” and “cited-by” papers.
+This improves the performance metrics, but comes at a cost of coverage metrics.
+To understand whether this reduction is caused by the RAG setup or the model, we include GPT 4o-mini in the same RAG setup, which shows a significant improvement in all metrics.
+Model Model type R-L BS JS CS CGT CLLM Prec. Recall F1
+T5 fine-tuned 19.92 87.81 10.82 31.79 35.48 29.59 0.29 0.31 0.30
+BART fine-tuned 19.43 87.67 10.68 31.91 33.71 30.10 0.30 0.31 0.31
+Pegasus fine-tuned 20.15 87.66 10.71 33.39 29.28 25.27 0.25 0.26 0.26
+Llama 3 zero-shot 25.66 88.30 14.69 40.4 61.38 39.04 0.39 0.50 0.44
+GPT-3.5 zero-shot 24.24 87.08 14.65 43.12 76.62 46.65 0.47 0.67 0.55
+GPT 4o-mini zero-shot 16.57 86.02 8.70 32.29 57.65 19.76 0.20 0.31 0.24
+GPT-3.5 + RAG zero-shot + RAG 30.21 90.88 19.47 45.37 39.99 44.66 0.42 0.40 0.41
+GPT 4o-mini + RAG zero-shot + RAG 23.17 87.33 12.99 39.29 67.13 45.67 0.57 0.45 0.51
+Table 3: Results of models in “Coverage” (Coverage of Ground Truth Limitation (CGT), LLM Generated Limitation (CLLM), Precision, Recall, and F1-score) and “performance” metrics – Rouge-x, BLEU, BertScore (BS), Jaccard (JS) and Cosine (CS) similarity on the ACL dataset.
+In all metrics, a higher score denotes a better performance.
+7.3 Self-reported & Peer-review Limitation
+The ground truth here consists of author-stated limitations and peer-review limitations extracted from NeurIPS papers.
+We hypothesize that the RAG approaches should be beneficial for this dataset, as the reviewers are more likely to point out limitations from external sources, such as cited in/by or semantically similar papers.
+Therefore, we use this dataset to compare different RAG approaches with GPT 4o-mini as the baseline LLM, as the previous experiments (Table 3) suggest that it has the highest propensity of improvement with RAG.
+RAG Index C_GT C_LLM F1 R-L BS CS JS
+Not applicable 67.34 63.81 0.65 15.30 86.66 33.43 8.69
+100 Random Papers 60.31 48.87 0.52 16.39 86.88 32.34 8.99
+Cited In 69.27 62.59 0.65 14.07 86.37 33.23 8.05
+Cited By 68.45 64.01 0.65 14.49 86.34 34.09 8.36
+Cited In + Cited By 68.84 64.87 0.67 14.35 86.35 33.94 8.28
+Cited In + By + Semantically Similar 5 Papers 68.02 63.38 0.65 14.59 86.39 33.72 8.38
+Table 4: Coverage evaluation of multiple types of RAG vector database settings in NeurIPS 21-22 dataset with GPT 4o-mini as the base LLM.
+Table 4 presents the performances with different RAG indices, with the first row representing vanilla LLM (no RAG).
+When the index is built with 100 random papers, the F1 score drops (-0.13) compared to the zero-shot approach.
+A combination of “cited in” and “cited by” papers achieves the highest F1 score of 0.67 – an increase of 0.02 over the baseline.
+However, when we further add the top five semantically related papers retrieved via the Semantic Scholar API, the F1 score reduces (-0.02), indicating that including loosely related content can introduce noise and reduce overall precision.
+However, the performance metrics present a somewhat different story.
+Semantically related sentences from 100 randomly selected papers yield the highest scores across multiple metrics, including ROUGE-L, BERTScore, BLEU, and Jaccard similarity.
+We believe this is due to the vector database containing diverse texts, which are not semantically or n-gram overlapping with the ground truth.
+This perhaps also shows the brittleness of performance metrics for evaluating the generation quality of limitations.
 7.4 Ablation Study
-We ablate in two dimensions a) the size of the input text. b) the ground truth
 a.
-Size of the input text: Table 5 presents results using GPT-4o mini.
-Moving from the top three sections to all sections as input improves every metric across all ground-truth types.
-The only exceptions are a few cases, such as BERTScore and LLM Coverage (CLLM) when using OpenReview as a ground truth.
-We also experimented ablation study with other models (Table 6), such as Mistral 7B and Llama 3 8B on NeurIPS data.
-Using all sections as input yields more detailed outputs, boosting LLM coverage by 8.35 points.
-However, it slightly reduces ground-truth coverage, CGT (-0.29), ROUGE-L (+0.64), and BERTScore (+0.36).
-Llama 3 8B shows the opposite trend: using all sections significantly degrades its performance.
-This likely reflects Llama’s limited context window, which hinders its ability to process longer inputs and generate coherent limitations.
+Size of the input text: We investigate the effect of the length of the input text on the generator models with an ablation study (Table 5).
+We use a) GPT-4o mini + RAG and b) Llama-3.1-8B, as these are the best-performing systems in the RAG and vanilla LLM setups, considering the author-mentioned limitations, the review-mentioned ones, and their combinations.
+When using GPT-4o-mini with RAG, expanding the context from the top-3 sections to all available sections generally increased pointwise scores for the author-written ground truth: CGT (+0.94), CLLM (+0.82), and F1 (+0.02).
+A similar trend was observed for the reviewer-suggested and combined ground truths, with the exception of a slight dip in the CGT score for the combined (Auth + Rev) case.
+By contrast, most NLP-based metrics (e.g., ROUGE, BERTScore, and cosine) slightly decreased with all-section inputs.
+Taken together, this indicates that using only the top-3 sections is a cost-effective alternative in this setup: minor drops in pointwise metrics, small gains (or less drop) in NLP metrics, and no large performance loss overall.
+In Llama-3.1-8B, however, we observe the opposite trend.
+Moving to all sections produces a large F1 gain (+0.13) for the combined ground truth and improves most NLP-based metrics.
+This suggests the smaller Llama-3.1-8B benefits from the full-paper context to generate higher-quality limitations, whereas truncating to the top-3 sections leaves it under-informed.
 b.
-Ground truth: Table 5 compares zero-shot GPT-4o mini results against three ground-truth sets: author-stated limitations, OpenReview comments, and their combination.
-Using both Limitations + OpenReview yields higher scores on nearly every metric compared to Limitations alone, except for CSKeyword, JSKeyword, and CGT.
-This demonstrates that reviewer feedback enriches the ground truth, improving overall performance, but because the combined ground truth is larger, its relative coverage (CGT) can decrease.
-We also experimented with Mistral 7b (Table 7, Appendix), incorporating OpenReivew increased CLLM, Cosine similarity, and F1 score but marginal drops in other metrics.
+Retriever Method: On the NeurIPS dataset, we evaluate our LLM re-ranker against a vanilla retriever baseline, both operating within a RAG framework with GPT-4o mini generator (Table 6).
+While the baseline simply retrieves the top 3 chunks using a FAISS+BM25 search, our method re-ranks the top 20 chunks, leading to substantial gains in CGT (+28.5), CLLM (+15.53), and the F1 score (+0.24).
 Metric Input Ground Truth
-Sec Auth OPR Auth + OPR
-R-L 3 16.57 18.55 17.07
-(+0.5)
-R-L All 17.11 (
+Sec Auth Rev. Auth + Rev
+GPT 4o-mini + RAG
+R-L 3 16.93 11.92 14.44
+All 16.73 (
+$\downarrow$
+12.12 (
+$\uparrow$
+14.35 (
+$\downarrow$
+BS 3 87.27 86.26 86.43
+All 87.15 (
+$\downarrow$
+86.21 (
+$\downarrow$
+86.35 (
+$\downarrow$
+CS 3 36.03 29.16 33.88
+All 35.52 (
+$\downarrow$
+29.81 (
 $\uparrow$
@@
 $\uparrow$
-18.83
-(+1.72)
-BS 3 86.02 86.50 86.25
-(+0.23)
-BS All 86.24 (
+CGT 3 82.60 61.19 69.78
+All 83.54 (
+$\uparrow$
+61.93 (
 $\uparrow$
@@
 $\downarrow$
-86.67
-(+0.43)
-JStext 3 8.70 10.17 9.12
-(+0.42)
-JStext All 9.22 (
+CLLM 3 29.83 62.69 61.59
+All 30.65 (
 $\uparrow$
@@
 $\uparrow$
-10.38
-(+1.16)
-CStext 3 32.29 38.35 35.42
-(+3.13)
-CStext All 33.23 (
-$\downarrow$
 64.87 (
-38.87
-(+5.64)
-CGT 3 57.65 54.79 45.69
-(-12.41)
-CGT All 61.55 (
+$\uparrow$
+F1 3 0.40 0.62 0.64
+All 0.42 (
 $\uparrow$
@@
 $\uparrow$
-49.43
-(-12.12)
-CLLM 3 19.76 56.79 76.27
-(+57.07)
-CLLM All 40.04 (
+0.67 (
+$\uparrow$
+Llama 3.1 8B
+R-L 3 17.67 12.30 15.03
+All 17.76 (
+$\uparrow$
+13.75 (
 $\uparrow$
@@
 $\downarrow$
-42.59
-(+2.55)
-F1 3 0.24 0.55 0.54
-(+0.3)
-F1 All 0.45 (
+BS 3 87.34 86.53 86.65
+All 87.54 (
 $\uparrow$
@@
 $\uparrow$
-0.44
-(-0.01)
-Table 5: GPT 4o mini results in “coverage” (Coverage of Ground Truth Limitation (CGT), LLM Generated Limitation (CLLM), F1-score) and “performance” (Rouge-x, BertScore (over entire text and just heading), Jaccard (JS) and Cosine (CS) similarity over text and heading) (NeurIPS data).
-Metric Input (3 Sec.) Input (All Sec.)
-Mistral 7B
-Rouge-L 14.59 13.95 (-0.64)
-BERTScore 84.59 84.23 (−0.36)
-JStext 7.70 7.49 (−0.21)
-CStext 30.95 31.82 (+0.87)
-BLEU 0.30 0.30 (0)
-CGT 38.28 37.99 (−0.29)
-CLLM 22.09 30.44 (+8.35)
-F1 score 0.24 0.32 (+0.08)
-Llama 3 8B
-Rouge-L 25.66 22.78 (−2.88)
-BERTScore 88.30 87.29 (−1.01)
-JStext 14.69 11.16 (−3.53)
-CStext 40.40 34.80 (−5.60)
-JSkeyword 20.28 20.31 (+0.03)
-CGT 61.38 18.01 (−43.37)
-CLLM 39.04 5.01 (−34.03)
-F1 score 0.44 0.05 (−0.39)
-Table 6: Ablation study comparing Mistral 7B and Llama 3 8B on NeurIPS data.
-We report metrics with only the top 3 sections versus all sections as input, using author-mentioned limitations as ground truth.
-8 Discussion
-Our results demonstrate that leveraging external OpenReview feedback alongside author-stated limitations substantially enriches the scope and quality of LLM-generated limitations, improving performance and coverage.
-This indicates the model not only reproduces author-reported issues but also uncovers additional shortcomings highlighted by peer reviewers.
-Similarly, integrating a Retrieval-Augmented Generation (RAG) module with GPT-3.5/4 improved every metric by grounding the LLM in relevant external context.
-The RAG-enhanced model produced more concise, targeted limitations (fewer sentences on average), trading a slight dip in recall with improved performance.
-Using all sections marginally improved performance then considering 3 section in most of the metrics.
-Moreover, our PointWise (PW) framework breaks each limitation into fine-grained points.
-It employs an LLM "judge" to evaluate alignment one by one, where traditional lexical metrics can’t capture depth and relevance.
-Finally, chunking long documents ensures no critical content is lost outside an LLM’s context window, and RAG helps focus on significant, relevant limitations outside the current paper.
-Moreover, models like GPT-4o mini deliver consistently strong performance with high correlation with humans as extractors and judges, whereas others (e.g., Llama) occasionally introduce noise.
-These findings affirm that combining external review data, targeted section selection, RAG augmentation, and a detailed PointWise evaluation yields more trustworthy and comprehensive LLM-generated limitations—and paves the way for future work in automated scientific limitations.
+87.06 (
+$\uparrow$
+CS 3 31.82 25.16 30.72
+All 31.99 (
+$\uparrow$
+27.47 (
+$\uparrow$
+29.99 (
+$\downarrow$
+CGT 3 63.52 42.79 44.74
+All 64.47 (
+$\uparrow$
+57.26 (
+$\uparrow$
+62.04 (
+$\uparrow$
+CLLM 3 33.32 44.67 48.93
+All 27.75 (
+$\downarrow$
+55.89 (
+$\uparrow$
+58.86 (
+$\uparrow$
+F1 3 0.39 0.43 0.46
+All 0.34 (
+$\downarrow$
+0.56 (
+$\uparrow$
+0.59 (
+$\uparrow$
+Table 5: Ablation study with GPT 4o-mini + RAG and Llama 3.1 8B results in “coverage” (Coverage of Ground Truth Limitation (CGT), LLM Generated Limitation (CLLM), F1-score) and “performance” (Rouge-x, BertScore (BS), and Cosine (CS) similarity in the NeurIPS data.
+Model VD C_GT C_LLM F1
+GPT 4o-mini Vanila k =3 40.34 49.34 0.43
+GPT 4o-mini LLM re-ranker 68.84 64.87 0.67
+Table 6: Performance between different retriever approaches in VD (Vector Database) in RAG (vanilla RAG (considering top 3 chunks) vs LLM re-ranker)
+Our findings demonstrate that a multi-faceted approach, combining curated external data with targeted retrieval, significantly enhances the generation of scientific limitations.
+This is especially evident when we use limitations extracted from reviews in the ground truth, as the use of “cited in” and “cited by” papers in the RAG index achieves the highest F1 score.
+We also observe that the length of the input to the generator model has a different effect in the vanilla LLM and RAG setup.
+It might be beneficial to use full paper texts for smaller models, but larger models in RAG setups can perform reasonably well with the most important parts of a paper.
 8 Conclusion
-We present a new approach for automatically extracting, generating and evaluating limitations in scientific articles.
-Our method explores chunk analysis, accommodating top sections of the entire paper and integrating OpenReview feedback to capture perspectives beyond those of the original authors.
-To evaluate the effectiveness of our system, we introduce a granular text evaluation framework that breaks down limitations into more minor points and employs LLMs as a Judge for assessing topic-level alignment.
-Combining standard NLP metrics with a keyword-based analysis provides a more nuanced and reliable assessment.
-Our findings are further validated by human review, which reveals a strong alignment between human judgments and the LLM-driven extraction process, including when the LLM is used as a judge.
+We present a new approach for automatically extracting, generating, and evaluating limitations in scientific articles.
+Our method explores incorporating cited works, accommodating top sections of the entire paper, and integrating review feedback to capture perspectives beyond those of the original authors.
+To evaluate the effectiveness of our system, we introduce a granular text evaluation framework that breaks down limitations into more minor points and employs LLMs as a Judge for assessing alignment.
+Human review validates our extraction and LLM-as-Judge pipeline, showing strong agreement with expert judgments.
 Limitations
@@
 Due to high API costs, we did not experiment with GPT-4 or GPT-4o; instead, we opted for GPT-4o Mini as a cost-effective alternative.
-Similarly, to avoid the high expense of maintaining an extensive vector database for RAG with closed LLMs, we limited our retrieval corpus to 100 randomly selected papers.
-We did not evaluate Open LLMs with RAG, such as Llama and Mistral, to assess their performance in retrieval-augmented settings.
 While we incorporated OpenReview comments for NeurIPS papers, we could not find them for ACL papers.
-Furthermore, we relied only on GPT-4o Mini as the evaluation judge and did not experiment with other LLMs for assessment.
+Furthermore, we relied on GPT-4o Mini as the evaluation judge.
 To evaluate the effectiveness of LLMs as both text extractors and judges, we conducted a human annotation study with 200 samples and only three annotators.
-We used GPT 4o mini for text extraction, text generation, and LLM as a judge.
-We experimented with other LLMs such as Llama but GPT 4o mini performs better.
-To rely on LLM for limitation extraction and evaluation ther are some potential biases.
-  • Self-Validation Bias: We used the same LLM for text generation and evaluation in experiments. Where a model can favor its style or logic.
-  • Positional bias: For LLM as a judge, we have multiple texts to check whether both are similar. Here, LLM can have positional bias, which is more centered on the first input text.
-  • Confirmation bias: We experimented with the same LLM (GPT 4o mini) for text extraction and generation.
-Even though both are used in an isolated environment with zero shot (no fine-tuning), the same LLM can produce confirmation bias on some common/generic limitations such as “Generalizability issue" or “small sample size" which reduces novelty.
-  • Hallucination and Overconfidence: When extracting limitations, LLM might sometimes invent limitations that aren’t grounded in the source text.
-To mitigate self-validation bias, we compared our LLM’s judgments against human evaluation, tested alternative models for generation and judgment (Llama and Mistral), and incorporated RAG.
-We changed the order and considered the consistent one to overcome the positional bias problem.
-Moreover, we integrated RAG with LLM for text generation, where LLM can get diverse evidence or force diversity sampling on generations to minimize the confirmation bias problem.
-To overcome hallucination and overconfidence, we evaluate the extraction process by LLM with three annotators to determine whether LLM extracts limitations without inventing anything new.
-However, these approaches are not sufficient.
-As part of our future work, we plan to expand the dataset to include papers from a broader range of domains, such as arXiv submissions from diverse subject areas (e.g., bioinformatics, cognitive science) and other peer-reviewed venues beyond ACL and NeurIPS.
-We also aim to incorporate more review texts from platforms, where available, across different conferences.
-This expansion will allow us to test the robustness of our limitation extraction models across domains and improve generalizability.
-We will explore broader contexts from open LLMs incorporating RAG, and integrating more extensive multi-agent LLM frameworks.
-Moreover, we will gather annotations from a larger and more diverse pool of experts, enabling a deeper analysis of generated limitations.
-In doing so, we aim to make our framework more robust and broadly applicable while maintaining its core strength of providing granular, interpretable assessments of LLM-generated limitations.
+A key threat to validity is contamination bias, when evaluation examples (or close paraphrases) appear in a model’s training data, artificially inflating performance.
+To guard against this, we tested whether GPT-4o mini had been trained on our NeurIPS 2021–2022 dataset by providing only each paper’s title and prompting it to summarize the content and identify limitations.
+In every case, the model replied with a disclaimer indicating unfamiliarity with the specific work (e.g., “I am not familiar with the specific paper titled …”).
+This consistent outcome suggests the model lacked prior exposure to the full texts, supporting the integrity of our evaluation.
+While we selected GPT-4o mini for text extraction, generation, and evaluation due to its superior performance, relying on a single LLM for these roles introduces several potential biases.
+We took specific steps to mitigate these risks: To counter self-validation bias, where the model might favor its own output, we cross-referenced its judgments with human evaluations and incorporated RAG.
+For positional bias, where the model may favor the first input when comparing texts, we swapped the input order to ensure consistent results.
+To reduce confirmation bias, the tendency to generate generic limitations, we used RAG to introduce more diverse evidence.
+Finally, to check for hallucinations, three human annotators verified that all extracted limitations were grounded in the source text.
+Although these strategies are crucial for improving reliability, we acknowledge that they do not completely eliminate these inherent biases.
+For future work, we will expand our dataset to more diverse domains (e.g., bioinformatics, cognitive science) to test the cross-domain robustness of our models.
+We also plan to enhance our generation framework by exploring more advanced multi-agent and open-source LLMs via RAG.
+Finally, we will scale our human validation efforts with a larger, more diverse pool of expert annotators to enable a deeper and more reliable analysis.
 Ethics Statement
-This research adheres to the ethical standards in scientific research.
-All data used in this study, including research papers and OpenReview feedback, were obtained from publicly available sources and used in compliance with their respective usage policies.
-No private or sensitive information was accessed or utilized.
-We acknowledge the potential biases inherent in large language models (LLMs) and have taken steps to mitigate their impact.
-For instance, we conducted human evaluations to validate the outputs of LLMs, ensuring alignment with human judgment and reducing the risk of generating misleading or inaccurate content.
-Additionally, our evaluation framework emphasizes transparency and granularity, enabling a more nuanced assessment of generated limitations.
-For the user study, we engaged three computer science graduate students; they are no longer involved in this research.
-Their participation was voluntary, and they had no conflicts of interest regarding the study.
-We are committed to promoting the responsible use of AI in scientific research.
-Our work aims to enhance scientific studies’ transparency, reliability, and reproducibility by generating and evaluating limitations systematically and ethically.
-We encourage further research to address potential biases in LLMs and improve their application in scientific contexts.
-Our study didn’t filter out any dataset based on race, ethnicity, color, location, gender, or other discriminatory attributes.
-Therefore, we are dedicated to complying with the ethical code and policy attributed to ACL.
-Large language models inherently carry biases and risks derived from their training corpora—confirmation bias toward “safe” limitations (e.g., small sample size), fluency bias favoring well-written over factually grounded outputs, verbosity bias rewarding longer responses, and self-validation bias when the same model is used for extraction, generation, and evaluation.
-To mitigate these, we ground all generations in source content and peer reviews via Retrieval-Augmented Generation (RAG), diversify our ground truth by incorporating human-authored OpenReview critiques, and conduct parallel human evaluations to detect overconfidence and evaluation bias.
-RAG-based outputs are demonstrably more concise, reducing verbosity bias, and our multi-model judge setup (including Llama and Mistral alongside GPT-4o mini) helps break circularity.
-We acknowledge that further work is needed to quantify these biases rigorously and plan to investigate cross-domain robustness and fairness in future studies.
+This research adheres to ACL ethical standards.
+All data, including research papers and OpenReview feedback, were sourced from public repositories in compliance with their usage policies and were not filtered based on discriminatory attributes.
+Our user study involved three computer science graduate students who participated voluntarily with no conflicts of interest.
+We acknowledge and address inherent LLM risks, including biases from training corpora, confirmation bias toward “safe” limitations, fluency and verbosity biases favoring longer or well-written outputs, and self-validation bias when using the same model for multiple tasks.
+To mitigate these, we (1) ground all generations in source content and peer reviews via a RAG framework to improve factuality and reduce verbosity; (2) diversify our ground truth by incorporating human-authored OpenReview critiques; (3) use multiple models to break self-validation circularity; and (4) conduct parallel human evaluations to detect overconfidence and other model-specific biases.
+We recognize that further work is needed to rigorously quantify these issues and plan to investigate cross-domain robustness in future studies.
 Appendix A Appendix
@@
 This setup enables us to examine how restricting the analysis to specific sections affects coverage and matching performance.
-We used three distinct prompts to check the topic level similarity between ground truth limitations and LLM-generated limitations (Figure 3, Appendix).
+We used three distinct prompts to check the topic-level similarity between ground truth limitations and LLM-generated limitations (Figure 5, Appendix).
 To overcome the position bias, we choose the consistent one.
-Metric Limitation Limitaiton + OPR
-Rouge-L 14.59 14.52 (-0.07)
-BERTScore 84.59 84.54 (-0.05)
-JStext 7.70 7.62 (-0.08)
-CStext 30.95 32.02 (+1.07)
-CGT 38.28 32.51 (-5.77)
-CLLM 22.09 44.46 (+22.37)
-F1 score 0.24 0.34 (+0.1)
-Table 7: Ablation study with Mistral 7b in NeurIPS data, considering Limitation and Limitation + OpenReview (OPR) as ground truth. Top 3 sections are input here.
-We also compared BERTScore between the Traditional approach vs the Pointwise approach, where the Pointwise approach preserves nuanced details and showed better performance (Table 8, Appendix).
-Approach Traditional way Pointwise approach
-T5 82.60 87.81 (+5.21)
-BART 83.09 87.67 (+ 4.58)
-Pegasus 82.76 87.66 (+4.9)
-GPT 3.5 83.79 87.08 (+3.29)
-GPT 4o mini 82.66 86.02 (+3.36)
-Table 8: Comparison of Traditional Approach and Pointwise Approach. BERTScore of each model in ACL data (Limitation is ground truth)
 GPT-4 Llama HE1 HE2 HE3
-GPT-4 1 0.71 0.9 0.92 0.95
+GPT-4o mini 1 0.71 0.9 0.92 0.95
 Llama - 1 0.81 0.79 0.76
@@
 Table 7: Evaluating how good LLM ‘as a Judge’ by checking Human Expert (HE) and model (GPT-4o mini, Llama-3.1 400B) agreement in determining whether an extracted limitation matches a generated one (in PointWise Evaluation).
-Model R-1 R-2 R-L BertScore BertScoreheading JStext CStext BLEU CSkeyword JSkeyword
-T5 27.13 5.11 19.92 87.81 - 10.82 31.79 0.4 18.69 11.26
-BART 26.05 5.49 19.43 87.67 - 10.68 31.91 0.7 19.89 12.19
-Pegasus 27.0 5.46 20.15 87.66 - 10.71 33.39 0.6 17.65 10.58
-GPT 3.5 32.74 10.92 24.24 87.08 85.50 14.65 43.12 3.54 26.54 17.46
-Llama 3 33.98 12.02 25.66 88.30 85.30 14.69 40.4 3.6 30.33 20.28
-GPT 4o-mini 22.90 3.99 16.57 86.02 75.02 8.70 32.29 0.33 16.20 9.85
-GPT 3.5 + RAG 38.11 16.31 30.21 90.88 87.65 19.47 45.37 6.64 35.74 24.65
-GPT 4o mini + RAG 31 9.95 23.17 87.33 - 12.99 39.29 2.69 26.47 17.19
-Table 10: Results of models in “performance” metrics – Rouge-x, BLEU, BertScore (over entire text and just heading), Jaccard (JS) and Cosine (CS) similarity over text and heading (ACL data).
-Metric ACL NeurIPS PeerJ
-Rouge-L 16.57 17.07 10.62
-BERTScore 86.02 86.25 83.26
-JStext 8.70 9.12 5.81
-CStext 32.29 35.42 26.19
-CSkeyword 16.20 15.67 2.89
-JSkeyword 9.85 9.31 1.66
-BLEU 0.33 0.4 0.01
-CGT 57.65 45.69 55.07
-CLLM 19.76 76.27 35.87
-F1 score 0.24 0.54 0.39
-Table 11: Comparison among ACL and NeurIPS PeerJ data in terms of Coverage and Performance.
-Full text is input, using GPT 4o mini to generate limitations, and Ground truth is Limitation + Peer Review (for NeurIPS and PeerJ), ACL has only author-mentioned Limitations.
-Metric Input Ground Truth
-Sec Auth OPR Auth + OPR
-R-1 3 22.90 26.13 23.69 ( +0.79)
-R-1 All 23.74 (
-$\uparrow$
-26.36 (
-$\uparrow$
-26.36 (+2.62)
-R-2 3 3.99 5.44 4.75
-(+0.76)
-R-2 All 4.51 (
-$\uparrow$
-5.41 (
-$\downarrow$
-5.64
-(+1.13)
-R-L 3 16.57 18.55 17.07
-(+0.5)
-R-L All 17.11 (
-$\uparrow$
-18.65 (
-$\uparrow$
-18.83
-(+1.72)
-BS 3 86.02 86.50 86.25
-(+0.23)
-BS All 86.24 (
-$\uparrow$
-85.96 (
-$\downarrow$
-86.67
-(+0.43)
-BSheading 3 75.02 85.79 69.50
-(-5.52)
-BSheading All 85.37 (
-$\uparrow$
-85.75 (
-$\downarrow$
-85.59
-(+0.22)
-JStext 3 8.70 10.17 9.12
-(+0.42)
-JStext All 9.22 (
-$\uparrow$
-10.23 (
-$\uparrow$
-10.38
-(+1.16)
-CStext 3 32.29 38.35 35.42
-(+3.13)
-CStext All 33.23 (
-$\downarrow$
-38.74 (
-38.87
-(+5.64)
-BLEU 3 0.33 0.5 0.4
-(+0.07)
-BLEU All 0.4 (
-$\uparrow$
-0.5 (
-$\uparrow$
-0.6
-(+0.2)
-CSKeyword 3 16.20 15.15 15.67
-(-0.53)
-CSKeyword All 18.14 (
-$\uparrow$
-15.42 (
-$\uparrow$
-16.83
-(-2.72)
-JSKeyword 3 9.85 8.99 9.31
-(-0.54)
-JSKeyword All 11.10 (
-$\uparrow$
-9.21 (
-$\uparrow$
-10.19
-(-0.91)
-CGT 3 57.65 54.79 45.69
-(-12.41)
-CGT All 61.55 (
-$\uparrow$
-56.55 (
-$\uparrow$
-49.43
-(-12.12)
-CLLM 3 19.76 56.79 76.27
-(+57.07)
-CLLM All 40.04 (
-$\uparrow$
-56.77 (
-$\downarrow$
-42.59
-(+2.55)
-Pr. 3 0.20 0.56 0.65
-(+0.46)
-Pr. All 0.4 (
-$\uparrow$
-0.57 (
-$\uparrow$
-0.43
-(+0.03)
-Rec. 3 0.31 0.55 0.46
-(+0.15)
-Rec. All 0.51 (
-$\uparrow$
-0.56 (
-$\uparrow$
-0.46
-(-0.05)
-F1 3 0.24 0.55 0.54
-(+0.3)
-F1 All 0.45 (
-$\uparrow$
-0.56 (
-$\uparrow$
-0.44
-(-0.01)
-Table 12: GPT 4o mini results in “coverage” (Coverage of Ground Truth Limitation (CGT), LLM Generated Limitation (CLLM), Precision, Recall, and F1-score) and “performance” (Rouge-x, BertScore (over entire text and just heading), Jaccard (JS) and Cosine (CS) similarity over text and heading) metrics when different input sizes and sources are used for limitation extraction.
-Increasing input length and including external reviews improve generation results (NeurIPS data).
-Metric GPT 4o mini GPT 4o mini + RAG
-Rouge-1 23.69 24.69 (+1.0)
-Rouge-2 4.75 5.24 (+0.49)
-Rouge-L 17.07 18.10 (+1.03)
-BERTScore 86.25 86.45 (+0.2)
-BSheading 69.50 74.74 (+5.24)
-JStext 9.12 9.72 (+0.6)
-CStext 35.42 34.71 (-0.71)
-CSkeyword 15.67 15.86 (+0.19)
-JSkeyword 9.31 9.60 (+0.29)
-BLEU 0.4 0.7 (+0.3)
-CGT 45.69 61.76 (+16.07)
-CLLM 76.27 59.18 (-17.09)
-Precision 0.65 0.60 (-0.05)
-Recall 0.46 0.59 (+0.13)
-F1 score 0.54 0.60 (+0.06)
-Table 13: Coverage and Performance measure using GPT 4o mini + RAG in NeurIPS data. Ground truth is Limitation + OpenReview, and input is the top 3 sections alongside the abstract.
-Metric Input (3 sections) Input (All Sections)
-Rouge-1 21.32 20.5 (-0.82)
-Rouge-2 3.91 4.0 (+0.09)
-Rouge-L 14.59 13.95 (-0.64)
-BERTScore 84.59 84.23 (-0.36)
-JStext 7.70 7.49 (-0.21)
-CStext 30.95 31.82 (+0.87)
-BLEU 0.3 0.3 (0)
-CSkeyword 16.43 17.29 (+0.86)
-JSkeyword 10.00 10.59 (+0.59)
-CGT 38.28 37.99 (-0.29)
-CLLM 22.09 30.44 (+8.35)
-F1 score 0.24 0.32 (+0.08)
-Table 14: Ablation study with Mistral 7b in Neurips data. Considering top 3 sections and all sections as input. Author mentioned limitation is ground truth here.
-Metric Input (3 sections) Input (All Sections)
-Rouge-1 33.98 29.57 (-4.41)
-Rouge-2 12.02 9.43 (-2.59)
-Rouge-L 25.66 22.78 (-2.88)
-BERTScore 88.30 87.29 (-1.01)
-JStext 14.69 11.16 (-3.53)
-CStext 40.40 34.80 (-5.6)
-BLEU 3.61 1.22 (-2.39)
-CSkeyword 30.33 30.73 (+0.4)
-JSkeyword 20.28 20.31 (+0.03)
-CGT 61.38 18.01 (-43.37)
-CLLM 39.04 5.01 (-34.03)
-F1 score 0.44 0.05 (-0.39)
-Table 15: Ablation study with Llama 3 8b in ACL data. Considering top 3 sections and All sections as input. Author mentioned limitation is ground truth here.
-Metric Limitation Limitaiton + OpenReview
-Rouge-1 21.32 21.02 (-0.3)
-Rouge-2 3.91 3.74 (-0.17)
-Rouge-L 14.59 14.52 (-0.07)
-BERTScore 84.59 84.54 (-0.05)
-JStext 7.70 7.62 (-0.08)
-CStext 30.95 32.02 (+1.07)
-BLEU 0.3 0.2 (-0.1)
-CSkeyword 16.43 14.69 (-1.74)
-JSkeyword 10.00 8.76 (-1.24)
-CGT 38.28 32.51 (-5.77)
-CLLM 22.09 44.46 (+22.37)
-F1 score 0.24 0.34 (+0.1)
-Table 16: Ablation study with Mistral 7b in NeurIPS data, considering Limitation and Limitation + OpenReview as ground truth. Top 3 sections are input here.
-Approach Limitation OpenReview Limitaiton + OpenReview
-Mistral (Traditional) 70.67 70.19 70.38
-Mistral (Pointwise) 84.23 (+13.56) 84.59 (+14.4) 84.54 (+14.16)
-Llama (Traditional) 67.55 67.17 67.50
-Llama (Pointwise) 87.29 (+19.74) - -
-Table 17: BERTScore Comparison Between PointWise (Proposed) and Traditional Approaches Across three different Ground Truths (Limitaiton, OpenReview, Limitation + OpenReview) in NeurIPS Data
-Agreement User 1 User 2 User 3
-User 1 1 0.75 0.38
-User 2 - 1 0.55
-User 3 - - 1
-Table 18: Inter annotator agreement amont human annotator for GPT 4o mini as Extractor.
 Section Cosine Similarity
@@
 Table 8: Cosine Similarity between each section and the Limitation section.
-Model Name Top 3 Sections All Sections
-CGT CLLM F1 Score CGT CLLM F1 Score
-Llama 3 61.38 39.04 0.44 18.01 (-43.37) 5.01 (-34.03) 0.05 (-0.39)
-GPT-4o Mini 57.65 19.76 0.24 61.55 (+3.9) 40.28 (+20.52) 0.46 (+0.22)
-Mistral 38.28 22.09 0.24 37.99 (-0.29) 30.44 (+8.35) 0.32 (+0.08)
-Table 20: Performance (Coverage) of various models considering the top three sections versus the entire paper.
-Note: CGT, CLLM denotes Coverage of Ground Truth Limitation, Coverage of LLM Generated Limitation.
-The ’+’ sign in parentheses indicates the performance gain achieved by using all sections instead of only the top three sections.
+Prompt = ''' Here is the text containing extracted limitations. Please identify and list each limitation, ensuring that each one addresses a distinct topic or point. '''
+Figure 3: Prompt to extract limitations from ground truth text.
+Prompt = ''' You are a helpful, respectful, and honest assistant for generating limitations or shortcomings of a research paper.
+I am providing ’Abstract’, ’Introduction’, ’Related Work’, ’Methodology’, ’Experiment and Results’, ’Conclusion’, and other sections of a scientific paper alongside the related cited papers texts.
+Generate limitations based on these texts.
+'''
+Figure 4: Prompt to generate limitations from Input and cited papers text.
 Prompt 1 = ''' A tuple contains (list1, list2).
@@
 Your answer should be "Yes" or "No" with explanation.
-Figure 3: LLM as a Judge for each limitation.
-Note: We use three distinct prompts to verify consistency.
-Ground Truth Limitation LLM Generated Limitation BS BS(h) Match
-Error-Prone Models:
-- Evaluation at the grade school mathematics level indicates that LLMs are still error-prone, but this does not provide insight into their effectiveness in tutoring other subjects.
-Keywords: tutoring, LLMs, mathematics, models, subjects
-Reliance on Large Language Models (LLMs): The research paper’s dependence on LLMs as Intelligent Tutoring Systems (ITS) for elementary school mathematics may limit the generalizability of the findings.
-Keywords: tutoring, mathematics, elementary, models, language
-87.88 81.08 Yes
-Errors in Code Generation: The presence of significant errors in code generation indicates that LLMs still face challenges in accurately performing tutoring tasks.
-Keywords: tutoring, LLMs, errors, code, challenges
-90.05 81.38 Yes
-Table 21: Ground Truth Limitation matched with multiple LLM Generated Text in PointWise Evaluation.
-Note: BS, BS(h), Match denotes BERTScore, BERTSCore (heading), and LLM as a judge decision.
-Prompt 1 = ''' Here is the text containing extracted limitations. Please identify and list each limitation, ensuring that each one addresses a distinct topic or point. '''
-Figure 4: Prompt to extract limitations from ground truth text.
-Ground Truth Limitation LLM Generated Limitation BS BS(h) Match
-Quality of Representations vs.
-Range of Words: The primary limitation is not the quality of the word representations produced, but rather the restricted range of words that the methodologies can be applied to.
-Keywords: representations, words, limitation, methodologies, range
-Dictionary Limitations: Dictionaries may not include definitions for all word forms in these languages, which restricts the applicability of the proposed approach.
-Keywords: dictionaries, dictionary, languages, forms, definitions
-89.12 83.08 Yes
-Dependence on Bilingual Dictionaries: The approach relies on bilingual dictionaries, which may not cover all necessary vocabulary or nuances of the languages.
-Keywords: bilingual, dictionaries, languages, vocabulary, relies
-91.71 82.78 Yes
-Table 22: LLM Generated Limitation matched with multiple Ground Truth Limitation.
-Note: BS, BS(h), Match denotes BERTScore, BERTSCore (heading), and LLM as a judge decision.
+Figure 5: LLM as a Judge for each limitation. We use three distinct prompts to verify consistency.
 Ground Truth Gen. Lim. GPT 4 Llama 3 User
-**Model Complexity Concerns**: - There is a question regarding whether the performance gains are due to increased model complexity rather than the proposed recursive mixing approach.
-**Potential for Increased Complexity**: Although the method is described as simple, the recursive nature of the approach may introduce complexity in implementation and understanding, particularly for practitioners who may not be familiar with the underlying concepts Yes Yes Yes
-**Computational Complexity**: The Dual-aspect Attention mechanism significantly increases computational complexity, which may affect the performance and practicality of the proposed method.
+Model Complexity Concerns: - There is a question regarding whether the performance gains are due to increased model complexity rather than the proposed recursive mixing approach.
+Potential for Increased Complexity: Although the method is described as simple, the recursive nature of the approach may introduce complexity in implementation and understanding, particularly for practitioners who may not be familiar with the underlying concepts Yes Yes Yes
+Computational Complexity: The Dual-aspect Attention mechanism significantly increases computational complexity, which may affect the performance and practicality of the proposed method.
 ’**Complexity of the Model**: The DACT arch, while innovative, introduces additional complexity compared to traditional models’, ’This complexity may lead to longer training times and increased computational resource requirements, which could be a barrier for practical applications in resource-constrained environments’ Yes Yes Yes
-**Insufficient Detailed Experimental Analysis**: The paper primarily provides quantitative results without a detailed experimental analysis, which limits the understanding of the findings.
-**Limited Discussion on Failure Cases**: While the paper discusses success and failure cases of existing methods, it may not provide sufficient insight into the specific conditions under which SCILL might fail, limiting the understanding of its robustness.
+Insufficient Detailed Experimental Analysis: The paper primarily provides quantitative results without a detailed experimental analysis, which limits the understanding of the findings.
+Limited Discussion on Failure Cases: While the paper discusses success and failure cases of existing methods, it may not provide sufficient insight into the specific conditions under which SCILL might fail, limiting the understanding of its robustness.
 Yes No Yes
-**Connection to Practical Methods**: There is an unknown connection between the theoretical findings and popular ensemble methods used in practice, which raises questions about the practical relevance of the work.
-**Unexplored Variants of Ensemble Methods**: The research does not explore other ensemble methods beyond those mentioned, such as boosting or stacking, which may have different theoretical properties and could provide additional insights into the interpolating regime No Yes Yes
-**Lack of Exploration of Alternative Methods**: The paper does not sufficiently explore or compare the performance of other methods like MFVI and SGLD in the context of covariate shift, which could provide a more comprehensive understanding of the problem.
-**Absence of Dataset Details**: The dataset section is marked as "nan," indicating a lack of information about the datasets used for experimentation’, ’This omission makes it difficult to assess the validity and applicability of the findings, as the choice of dataset can significantly influence the results No No No
-Table 23: LLM as a Judge (GPT 4, Llama 3) and User to decide whether Ground Truth Limitation and LLM Generated Limitation have the same topic or not.
-Note: Here GPT 4, Llama 3, and User decision regarding both limitation match or not.
+Connection to Practical Methods: There is an unknown connection between the theoretical findings and popular ensemble methods used in practice, which raises questions about the practical relevance of the work.
+Unexplored Variants of Ensemble Methods: The research does not explore other ensemble methods beyond those mentioned, such as boosting or stacking, which may have different theoretical properties and could provide additional insights into the interpolating regime No Yes Yes
+Lack of Exploration of Alternative Methods: The paper does not sufficiently explore or compare the performance of other methods like MFVI and SGLD in the context of covariate shift, which could provide a more comprehensive understanding of the problem.
+Absence of Dataset Details: The dataset section is marked as "nan," indicating a lack of information about the datasets used for experimentation’, ’This omission makes it difficult to assess the validity and applicability of the findings, as the choice of dataset can significantly influence the results No No No
+Table 9: Examples of Annotator, GPT 4o-mini, and LLama judgement on whether a generated limitation should be matched with a ground-truth limitation or not.
 Ground Truth Limitations: Our goal is to provide comprehensive study of learning algorithms, including their hyperparameters.
@@
 Specifically, we did not study how the type and configurations of the optimizer and the batch size would change the results; instead, we followed common setups in previous studies.
-'''   LLM Refined Limitations: ’1.
-Here are the limitations extracted from the provided limitation:’, ’2.
-.
-**Experimental Complexity**: The study is limited in the variety of design choices to reduce experimental complexity.’, ’3.
-.
-**Dataset Complexity**: The study utilizes relatively simple synthetic and realistic image datasets, which have a small number of generative factors.’, ’4.
-.
-**Generative Factors**: Each generative factor in the datasets follows a uniform distribution, limiting the diversity of the data.’, ’5.
-.
-**Model arch**: The focus is on studying the inductive bias related to the representation format while keeping the model arch design fixed, which may influence the results.’, ’ 6.
-.
-**Hyperparameter Scope**: The study does not explore hyperparameters beyond those related to latent representations.’, ’7.
-.
-**Optimizer and Batch Size**: The impact of different types and configurations of optimizers and batch sizes on the results is not examined, as the study adheres to common setups from previous research.’
+'''  LLM Extracted Limitations: 1.
+Here are the limitations extracted from the provided limitation:, 2.
+**Experimental Complexity**: The study is limited in the variety of design choices to reduce experimental complexity., 3.
+**Dataset Complexity**: The study utilizes relatively simple synthetic and realistic image datasets, which have a small number of generative factors., 4.
+**Generative Factors**: Each generative factor in the datasets follows a uniform distribution, limiting the diversity of the data., 5.
+**Model arch**: The focus is on studying the inductive bias related to the representation format while keeping the model arch design fixed, which may influence the results., 6.
+**Hyperparameter Scope**: The study does not explore hyperparameters beyond those related to latent representations., 7.
+**Optimizer and Batch Size**: The impact of different types and configurations of optimizers and batch sizes on the results is not examined, as the study adheres to common setups from previous research.
 Figure 6: Ground Truth Limitations and LLM Extracted Limitations in NeurIPS dataset.
@@
 Our work aims to establish preliminary evaluation of the humanmodel agreement and model annotations,and further validation of the automatic labels is necessary.
-'''   LLM Refined Limitations: Here are the limitations extracted from the provided limitation: 1.
+'''  LLM Extracted Limitations: Here are the limitations extracted from the provided limitation: 1.
 **Comments May Not Reflect Real Student Feedback**: The dataset consists of comments from users who have chosen to post publicly on YouTube, which may not accurately represent genuine student feedback.
@@
 Figure 7: Ground Truth Limitations and LLM Extracted Limitations in ACL dataset.
-Ground Truth OpenReview: Reviewer 1: This paper tries to tackle the domain generalization problem by finding flat minima.
+OpenReview: Reviewer 1: This paper tries to tackle the domain generalization problem by finding flat minima.
 It is interesting to investigate the loss landscape of the DNNs on domain generalization problems.
@@
 The proposed algorithm is relatively simple to implement in practice and perform well experimentally.
-Cons: My major concern is wrt to the novelty of the work and the connection/motivation between the proposed algorithm SWAD, domain generalization and the derived theoretical insights… '''   LLM Refined OpenReview: "1.
+Cons: My major concern is wrt to the novelty of the work and the connection/motivation between the proposed algorithm SWAD, domain generalization and the derived theoretical insights… '''  LLM extracted OpenReview: "1.
 .
@@
 **Potential for Broader Applications**: The method’s applicability beyond domain generalization is not explored, which could limit its significance and impact in the broader machine learning community.
-Figure 7: Ground Truth and LLM Extracted Limitations in OpenReview.
+Figure 8: Tool extracted OpenReview and LLM Extracted OpenReview.